### v2

# arXiv:2602.23189v2[math.AP] 02 Mar 2026

Abstract
We analyse a bi-fluid isentropic compressible Navier-Stokes system with barotropic pressure laws in a two-phase framework with equal pressure and single velocity. We focus on the rigorous analysis of the low Mach number limit under well-prepared initial data. Our main result shows that, as the Mach number tends to zero, the partial densities converge to constant states while the velocity field converges to a divergence-free vector field, and we recover the incompressible non-homogenous fluid system. The volume fractions remain nontrivial and are transported by the limit flow.
Our method is based on the introduction of suitable modulated quantities and a novel relative entropy functional adapted to the two-phase structure. The key novelty lies in an original comparison of $L^{1}$ and $L^{2}$ norms of the partial densities, exploiting the special structure of the bi-fluid system, together with a variant of Csiszar-Kullback-Pinsker inequality. It allows us to rigorously justify the convergence of weak solutions of the compressible system toward solutions of the corresponding incompressible limit system in the low Mach number regime.
1 Introduction
Compressible multiphase flow models are a fundamental class of nonlinear systems that arise in fluid mechanics and engineering applications. In many formulations, each phase satisfies its own balance laws, and additional closure relations are required to determine thermodynamic quantities and interfacial effects. Classical two-fluid models developed by Ishii and Hibiki [22], as well as the hyperbolic model introduced by Baer and Nunziato [1], exemplify this framework. Depending on the modeling assumptions, closures may be differential, introducing additional evolution equations, or algebraic, in which the pressure is expressed as a function of partial densities, volume fractions, or other state variables. In particular, algebraic isentropic closures induce a nontrivial coupling between the phases, which has important implications for the analytical structure of the system. From a mathematical perspective, these models have been studied in the compressible PDE community; see, for instance, the handbook chapter by DD. Bresch, B. Desjardins, J.M. Ghidaglia, E. Grenier and M. Hillairet [3], and the references therein. These works provide a mathematical formulation of the models, establish their main structural properties, and contain some of the first analytical results, marking the emergence of sustained interest in such systems within the theory of compressible PDEs.
The growing interest in compressible two-fluid flows with viscosity and weak-regularity data has led to significant mathematical advances, particularly in the viscous framework. Fundamental results on the existence of finite-energy weak solutions were established by Bresch, Mucha, and Zatorska [6] as well as by Vasseur, Wen, and Yu [45], highlighting suitable dissipation mechanisms. In the same spirit, Novotný and Pokorný [39] and Novotný [38] developed a weak solution theory of compressible bi-fluid models, while Jin, Kwon, Nečasová, and Novotný [24] introduced dissipative turbulent solutions and investigated their stability. Weak-strong uniqueness properties were further analyzed by Jin and Novotný [25] and by Li and Zatorska [35]. Moreover, Li and Zatorska [34] addressed weak solutions to the compressible inviscid two-fluid model, emphasizing the additional analytical difficulties arising in the absence of viscosity, whereas Piasecki and Zatorska [41] obtained maximal regularity results in the viscous setting. The large-time behavior of solutions with large initial data was studied by Li, Sun, and Zatorska [33]. Finally, the recent work of Li, Wang, and Zhang [31] established the non-existence of classical solutions in the presence of vacuum, further illustrating the analytical complexity of such models. Note that recently a preprint by Li, Lukáčová-Medvid’ová, Pokornỳ and Zatorska (see [32]) related to weak-strong uniqueness for two-phase systems has been posted on ArXiv independently of ours. These contributions, among others, form a coherent and evolving theory for compressible two-fluid flows.
In this work, we consider a compressible isentropic two-phase model with an algebraic closure and a single common velocity field for both phases. This assumption simplifies the structure of the system while retaining the essential coupling induced by the pressure law. After nondimensionalization, the Mach number $\varepsilon$ appears explicitly in front of the pressure gradient, producing a singular perturbation as $\varepsilon\to 0$. We are interested in the rigorous justification of the corresponding low Mach number limit under well-prepared initial data.
For single-phase flows, the low Mach number limit is by now well understood. For the compressible Euler and Navier–Stokes equations, convergence toward incompressible dynamics under well-prepared initial data has been established in various functional settings; see, for instance, the works of Klainerman and Majda [27], and subsequent developments in the context of symmetric hyperbolic systems and viscous flows.
Without aiming at exhaustiveness, we mention the following representative contributions. In the framework of weak solutions, the low Mach number limit has been justified in [36], [15], [14], and [4]. For strong solutions, one may refer to [21], [37], and [11]. A concise overview of these developments can be found in [19]. A comprehensive treatment of the compressible Navier–Stokes system with constant viscosities, with particular emphasis on the relative entropy method as a stability and convergence tool, is given in [18]. For more complex systems, such as compressible Navier–Stokes equations with density-dependent viscosities, the notion of relative entropy had to be revisited and adapted; see, for instance, [7] and [5]. We also mention works developing the relative entropy framework in related contexts, such as [17] and [43].
In contrast, significantly fewer results are available for multiphase flow models, even in simplified configurations. Only recently has attention been devoted to the formal derivation of low Mach number limits for various two-phase systems, including models with algebraic or PDE closures, as well as formulations involving one or two velocities. In these works, asymptotic expansions are typically performed as the Mach number tends to zero; see, for instance, [44] and [29]. We may also refer to Fanelli et al [16] for related incompressible limits in a reduced compressible MHD system.
In the present paper, in the two-phase setting with equal pressure, we introduce several relevant modulated quantities and propose a relative entropy functional based on a suitable combination of them. This approach requires the development of appropriate comparison tools in order to strengthen certain convergence properties.
Our main result establishes that, in the low Mach limit and under well-prepared initial data, the partial densities converge to constant values, while the velocity field converges to a divergence-free vector field. The volume fractions remain variable and are transported by the limit flow, so the mixture density evolves through advection of the phase fractions. This leads to a variable-density incompressible system in which the mixture density depends only on the transported volume fractions.
1.1 Bi-fluid compressible system with algebraic closure
In the present paper, we analyse a bi-fluid isentropic compressible Navier-Stokes system for barotropic pressures. More precisely, denoting the phases by ++ and -, we are considering the following system.
$\left\{\begin{aligned} &\partial_{t}(\alpha_{+}\rho_{+})+\operatorname{div}(\alpha_{+}\rho_{+}u)=0,\\ &\partial_{t}(\alpha_{-}\rho_{-})+\operatorname{div}(\alpha_{-}\rho_{-}u)=0,\\ &\partial_{t}((\alpha_{+}\rho_{+}+\alpha_{-}\rho_{-})u)+\operatorname{div}((\alpha_{+}\rho_{+}+\alpha_{-}\rho_{-})u\otimes u)+\nabla(\alpha_{+}p_{+}(\rho_{+})+\alpha_{-}p_{-}(\rho_{-}))\\ &\quad=\mu\Delta u+(\mu+\lambda)\nabla\operatorname{div}u,\\ &p_{i}(\rho_{i})=(\rho_{i})^{\gamma_{i}},\qquad\rho_{i}\geqslant 0,\qquad i=+,-,\\ &p_{+}(\rho_{+})=p_{-}(\rho_{-}),\\ &\alpha_{+}+\alpha_{-}=1,\qquad 0\leqslant\alpha_{+}\leqslant 1,\end{aligned}\right$. (1.1)
where $\rho_{i}$, $p_{i}$, $\alpha_{i}$ the density, pressure and volume fraction of phase $i$. The real numbers $\mu$ and $\lambda$ are the constant shear and bulk viscosities verifying, as usual,
$\mu>0\quad\text{and}\quad\lambda+2\mu>0.$
Let us notice that, thanks to the algebraic closure, the pressure term in the momentum equation can be simplified:
$\nabla(\alpha_{+}p_{+}(\rho_{+})+\alpha_{-}p_{-}(\rho_{-}))=\nabla(\alpha_{+}p_{+}(\rho_{+})+\alpha_{-}p_{+}(\rho_{+}))=\nabla p_{+}(\rho_{+}).$
In this system, we restrict ourselves to the case of a common velocity field, denoted by $u$. Physically, this corresponds to a regime of rapid mechanical relaxation between the phases, where pressure equilibrium is quickly achieved. Mathematically, such models can be rigorously derived as asymptotic limits of more general two-velocities, two-pressures systems, using homogenization (see, e.g., Bresch et al. [2]) or pressure-relaxation arguments (see [9], [8]). This assumption considerably simplifies the structure of the equations and is standard in the literature, see for instance [39]. Allowing for two distinct velocities would lead to a substantially more intricate coupling mechanism and raises major analytical difficulties; the rigorous study of such models remains a challenging and largely open problem.
The system is endowed with the Dirichlet boundary condition
$u(t,x)\lvert_{\partial\Omega}=0.$ (1.2)
on $(0,T)\times\partial\Omega$, where $\Omega$ is a sufficiently smooth bounded domain of $\mathbb{R}^{3}$ and $T>0$ is arbitrary large.
Remark 1.1.
The analysis can also be extended to periodic domains or to two-dimensional settings, possibly under additional assumptions on the initial data.
Let us denote by $\varepsilon$ the Mach number. The incompressible limit is the limit where the speed of the fluid becomes negligible compared to the speed of sound. Thus we reset the flow speed to the scale by replacing $u$ with $\varepsilon u^{\varepsilon}$. A particle will move a distance of order $1/\varepsilon$ during a time of order $1/\varepsilon^{2}$, meaning that the relevant time scale is $1/\varepsilon^{2}$. This is why we can use a change of variable like $t^{\varepsilon}=\varepsilon^{2}t$ and $x^{\varepsilon}=\varepsilon x$, with $t^{\varepsilon}$ and $x^{\varepsilon}$ of order $1$ (see [12]). Thus, we scale $u$ and $\rho$ in the following way
$u(t,x)=\varepsilon u^{\varepsilon}(t^{\varepsilon},x^{\varepsilon}),\quad\rho(t,x)=\rho^{\varepsilon}(t^{\varepsilon},x^{\varepsilon}).$
We can rewrite System (LABEL:system) in the following form:
$\left\{\begin{aligned} &\partial_{t}(\alpha_{+}^{\varepsilon}\rho_{+}^{\varepsilon})+\operatorname{div}(\alpha_{+}^{\varepsilon}\rho_{+}^{\varepsilon}u^{\varepsilon})=0,\\ &\partial_{t}(\alpha_{-}^{\varepsilon}\rho_{-}^{\varepsilon})+\operatorname{div}(\alpha_{-}^{\varepsilon}\rho_{-}^{\varepsilon}u^{\varepsilon})=0,\\ &\partial_{t}((\alpha_{+}^{\varepsilon}\rho_{+}^{\varepsilon}+\alpha_{-}^{\varepsilon}\rho_{-}^{\varepsilon})u^{\varepsilon})+\operatorname{div}((\alpha_{+}^{\varepsilon}\rho_{+}^{\varepsilon}+\alpha_{-}^{\varepsilon}\rho_{-}^{\varepsilon})u^{\varepsilon}\otimes u^{\varepsilon})+\boldsymbol{\frac{1}{\varepsilon^{2}}}\nabla p_{+}(\rho_{+}^{\varepsilon})\\ &\hskip 199.16928pt=\mu\Delta u^{\varepsilon}+(\mu+\lambda)\nabla\operatorname{div}u^{\varepsilon},\\ &p_{i}(\rho_{i}^{\varepsilon})=(\rho_{i}^{\varepsilon})^{\gamma_{i}},\qquad\rho_{i}^{\varepsilon}\geqslant 0,\qquad i=+,-\\ &p_{+}(\rho_{+}^{\varepsilon})=p_{-}(\rho_{-}^{\varepsilon}),\\ &\alpha_{+}^{\varepsilon}+\alpha_{-}^{\varepsilon}=1,\qquad 0\leqslant\alpha_{+}^{\varepsilon}\leqslant 1,\end{aligned}\right$. (1.3)
still considering the Dirichlet boundary condition (1.2).
As usual in the mathematical studies of bi-fluid models, we introduce the fractional masses $R_{\pm}^{\varepsilon}:=\alpha_{\pm}^{\varepsilon}\rho_{\pm}^{\varepsilon}$, the mixture density by $\rho^{\varepsilon}:=R_{+}^{\varepsilon}+R_{-}^{\varepsilon}$ and the momentum by $m^{\varepsilon}:=\rho^{\varepsilon}u^{\varepsilon}$.
Remark 1.2.
For $0\leqslant R_{+}^{\varepsilon}\leqslant\rho_{+}^{\varepsilon}$ and $m^{\varepsilon}$ given, if $\rho_{+}^{\varepsilon}>0$, we may define $\alpha_{+}^{\varepsilon}=R_{+}^{\varepsilon}/\rho_{+}^{\varepsilon}\in[0,1]$. Otherwise, we set $\alpha_{+}=1/2$ on $\{\rho_{+}^{\varepsilon}=0\}$. Next we define $\alpha_{-}^{\varepsilon}=1-\alpha_{+}^{\varepsilon}\in[0,1]$, $\rho_{-}^{\varepsilon}=(\rho_{+}^{\varepsilon})^{\gamma_{-}/\gamma_{+}}\geqslant 0$ (by the equality of pressures), $R_{-}^{\varepsilon}=\alpha_{-}^{\varepsilon}\rho_{-}^{\varepsilon}$ and $\rho^{\varepsilon}=R_{+}^{\varepsilon}+R_{-}^{\varepsilon}$. If $\rho_{+}^{\varepsilon}>0$, we get that $\rho_{-}^{\varepsilon},\rho^{\varepsilon}>0$ and we may define the velocity $u^{\varepsilon}=m^{\varepsilon}/\rho^{\varepsilon}$. Otherwise, we set $u^{\varepsilon}=0$ on $\{\rho_{+}^{\varepsilon}=0\}$.
For any functions $(\rho_{+,0}^{\varepsilon},R_{+,0}^{\varepsilon},m_{0}^{\varepsilon})\in L^{\gamma_{+}}\times L^{\gamma_{+}}\times L^{1}$ satisfying the following compatibility conditions
$0\leqslant R_{+,0}^{\varepsilon}\leqslant\rho_{+,0}^{\varepsilon}\text{\ a.e.},\quad\sqrt{\rho_{0}^{\varepsilon}}u_{0}^{\varepsilon}\in L^{2}$ (1.4)
where we have defined by the previous remark $\alpha_{\pm,0}^{\varepsilon}\in[0,1]$, $\rho_{-,0}^{\varepsilon}\geqslant 0$, $\rho^{\varepsilon}_{0}\geqslant 0$ and $u_{0}^{\varepsilon}$, we consider a global weak solution “à la Leray”, see Definition 2.1 for the precise meaning.
System (LABEL:system_eps) is a particular case of the system studied by [39] for which they obtain the existence of global-in-time finite-energy weak solutions for large initial data, but adding some conditions on $\gamma_{\pm}$. Note that [6] also gives a proof of the existence of global-in-time weak solutions for a semi-stationary bifluid system when $\gamma_{+}\neq\gamma_{-}$.
1.2 Bi-fluid incompressible system
Let us now investigate the Mach-number limit. As shown in [29], when we let the Mach number go to zero in (LABEL:system_eps), we obtain formally the following non-homogeneous viscous incompressible Navier-Stokes equations:
$\begin{cases}\rho_{+}=C_{0}^{\frac{1}{\gamma_{+}}},\qquad\rho_{-}=C_{0}^{\frac{1}{\gamma_{-}}},\\ \operatorname{div}u=0,\\ \partial_{t}\alpha_{+}+(u\cdot\nabla)\alpha_{+}=0,\\ \left(\rho_{+}\alpha_{+}+\rho_{-}\alpha_{-}\right)\left(\partial_{t}u+(u\cdot\nabla)u\right)+\nabla\Pi=\mu\Delta u,\\ \alpha_{+}+\alpha_{-}=1,\end{cases}$ (1.5)
where $C_{0}>0$ is a real number and $u$ satisfies the Dirichlet boundary condition (1.2). Let us notice that the first line implies the pressure equality:
$p_{+}(\rho_{+})=p_{-}(\rho_{-}).$
The densities $\rho_{\pm}$ are now constant, and the continuity equations on $(\alpha_{\pm}^{\varepsilon}\rho_{\pm}^{\varepsilon})$ became a transport equation on $\alpha_{\pm}$ by a divergence free vector field. It is then clear that the initial assumption $\alpha_{\pm,0}\in[0,1]$ remains true for all time $t>0$.
Denoting again the mixture density by $\rho:=\rho_{+}\alpha_{+}+\rho_{-}\alpha_{-}$ we get the classical non-homogeneous viscous incompressible Navier-Stokes equations where $\rho$ satisfies the transport equation by $u$.
In the case of nonhomogeneous incompressible flows with constant viscosities, the existence of weak and strong solutions has been investigated in several works. The existence of weak solutions was first established by Kazhikhov [26]. Concerning strong solutions, Ladyzhenskaya and Solonnikov [28] proved local existence, together with global existence for small initial data (and global solvability in two space dimensions). In sufficiently smooth domains, Lemoine [30] obtained local strong regular solutions, which are global for small regular data. Okamoto [40] proved existence and uniqueness of local weak solutions and obtained global solutions in three dimensions under smallness assumptions on the initial data. Simon [42] established the existence of global-in-time weak solutions under suitable weak assumptions on the initial data, and proved global strong solvability provided the initial density is not too small (and sufficiently regular). The solvability of nonstationary problems for nonhomogeneous incompressible fluids, as well as convergence in the vanishing viscosity limit, was studied by Itoh [23]. Choe and Kim [10] established local existence and uniqueness of strong solutions in three dimensions, including situations where the initial density vanishes on an open subset of the domain. More recently, Danchin [13] developed a well-posedness theory in critical Besov spaces and proved the existence of global-in-time unique solutions under suitable smallness assumptions on the initial data.
Among the vast existing literature, let us state a result of local existence of strong solutions of non-homogeneous viscous incompressible Navier-Stokes equations.
Theorem 1.3.
Let $(\rho_{0},u_{0})$ be some initial data such that $u_{0}\in H^{1}_{0}(\Omega)\cap H^{s}(\Omega)$, $\operatorname{div}u_{0}=0$ in $\Omega$ and $\rho_{0}\in L^{\infty}(\Omega)$, with $\overline{C}\geqslant\rho_{0}\geqslant\underline{C}>0$ for some constants $\overline{C},\underline{C}$. Then there exists $T>0$ and a (unique) solution $(\rho,u,\Pi)$ of (1.5) such taht $0<\underline{C}\leqslant\rho\leqslant\overline{C}$ that satisfy
$u\in L^{\infty}(0,T;H_{0}^{1}(\Omega)\cap H^{s}(\Omega)),\qquad\frac{\partial u}{\partial t}\in L^{2}(0,T;H_{0}^{1}(\Omega)),$
$\nabla^{s}u\in L^{2}(0,T;L^{6}(\Omega)),\qquad\nabla^{s+1}u\in L^{2}(0,T;L^{3/2}(\Omega)),$
$\rho\in L^{\infty}(0,T;H_{0}^{1}(\Omega)),$
$\frac{\partial\rho}{\partial t}\in L^{\infty}(0,T;L^{2}(\Omega)),\qquad\sqrt{\rho}u\in L^{\infty}(0,T;L^{2}(\Omega)),$
$\nabla\Pi\in L^{\infty}(0,T;L^{2}(\Omega))\cap L^{2}(0,T;L^{6}(\Omega)\cap W^{s-1,3/2}(\Omega)).$
Remark 1.4.
When we consider $(\rho,u)$ a regular solution to non-homogeneous viscous incompressible Navier-Stokes equations, it is clear that we may construct a solution of (1.5). Indeed, solving $\rho=\rho_{+}\alpha_{+}+\rho_{-}\alpha_{-}$ and $1=\alpha_{+}+\alpha_{-}$, we simply define
$\alpha_{+}=\frac{\rho-\rho_{-}}{\rho_{+}-\rho_{-}}\quad\text{and}\quad\alpha_{-}=-\frac{\rho-\rho_{+}}{\rho_{+}-\rho_{-}}$
which is well defined for $\gamma_{+}\neq\gamma_{-}$, as $C_{0}>0$. We observe that $\alpha_{\pm}$ satisfy the transport equation. Assuming initially $\alpha_{\pm,0}\in[\underline{a},\bar{a}]\subset(0,1)$, such a property then remains true. Let us notice that we can express $\underline{C}$ and $\overline{C}$ in terms of $\underline{a},\overline{a},C_{0},\gamma_{\pm}$.
Remark 1.5.
We will consider $s=3$. The existence result stated above is not optimal in terms of regularity for our subsequent arguments. For our purposes, a lower level of regularity would suffice, but this is not relevant here.
1.3 Main theorem
We can now state the main result of this paper, where we denote the fractional masses of the limit system by $R_{\pm 0}=\alpha_{\pm,0}\rho_{\pm,0}$.
Theorem 1.6 (Singular limit of the compressible two-phase system).
Let $C>0$, $\gamma_{\pm}\geqslant 2$ be three real numbers and $\alpha_{\pm,0}$ be two functions such that $\gamma_{+}\neq\gamma_{-}$ and $0<\underline{a}\leqslant\alpha_{\pm,0}\leqslant\overline{a}<1$. Let us assume that $(\alpha_{\pm,0},u_{0})$ are enough regular such that we can apply Theorem 1.3 for the initial data $(\rho_{0},u_{0})$, and let $T>0$ be the existence time of strong solution to System (1.5).
For each $\varepsilon>0$, let $(\rho_{+,0}^{\varepsilon},R_{+,0}^{\varepsilon},m_{0}^{\varepsilon})\in L^{\gamma_{+}}\times L^{\gamma_{+}}\times L^{1}$ such that compatibility conditions (1.4) are verified and the total masses are consistent with the limit:
$\int_{\Omega}R_{\pm,0}^{\varepsilon}\,dx=\int_{\Omega}R_{\pm,0}\,dx>0.$ (1.6)
Assume that
  1. (i)
    the initial energy is uniformly bounded, namely
    $\int_{\Omega}\frac{1}{2}\Big(\rho^{\varepsilon}_{0}|u^{\varepsilon}_{0}|^{2}\Big)(x)\,\mathrm{d}x+\int_{\Omega}\frac{\alpha_{+,0}^{\varepsilon}\rho_{+,0}^{\varepsilon\gamma_{+}}(x)}{\varepsilon^{2}(\gamma_{+}-1)}\,\mathrm{d}x+\int_{\Omega}\frac{\alpha_{-,0}^{\varepsilon}\rho_{-,0}^{\varepsilon\gamma_{-}}(x)}{\varepsilon^{2}(\gamma_{-}-1)}\,\mathrm{d}x\leqslant C,$ (1.7)
    and
    $\int_{\Omega}R_{+,0}^{\varepsilon}\ln_{+}\frac{R_{+,0}^{\varepsilon}}{R_{+,0}}+\int_{\Omega}R_{-,0}^{\varepsilon}\ln_{+}\frac{R_{-,0}^{\varepsilon}}{R_{-,0}}\leqslant C,$ (1.8)
    where, here and below, $C$ denotes positive constants independent of $\varepsilon$,
  2. (ii)
    the initial data is well prepared, in the sense
    $\begin{split}&\int_{\Omega}\frac{1}{2}\big(R^{\varepsilon}_{+,0}+R^{\varepsilon}_{-,0}\big)|u_{0}^{\varepsilon}-u_{0}|^{2}\,dx\\ &\quad+\int_{\Omega}\frac{\alpha_{+,0}^{\varepsilon}}{\varepsilon^{2}(\gamma_{+}-1)}\Big((\rho_{+,0}^{\varepsilon})^{\gamma_{+}}-\rho_{+,0}^{\gamma_{+}}-\gamma_{+}\rho_{+,0}^{\gamma_{+}-1}(\rho_{+,0}^{\varepsilon}-\rho_{+,0})\Big)\,dx\\ &\quad+\int_{\Omega}\frac{\alpha_{-,0}^{\varepsilon}}{\varepsilon^{2}(\gamma_{-}-1)}\Big((\rho_{-,0}^{\varepsilon})^{\gamma_{-}}-\rho_{-,0}^{\gamma_{-}}-\gamma_{-}\rho_{-,0}^{\gamma_{-}-1}(\rho_{-,0}^{\varepsilon}-\rho_{-,0})\Big)\,dx\longrightarrow 0.\end{split}$ (1.9)
    as $\varepsilon\to 0$.
Then, the corresponding weak solution $(\rho_{\pm}^{\varepsilon},\alpha_{\pm}^{\varepsilon},u^{\varepsilon})$ of the compressible two-phase Navier–Stokes system (LABEL:system_eps) on $(0,T)$, in the sense of Definition 2.1, converge strongly to the incompressible limit as $\varepsilon\to 0$, namely
$R_{\pm}^{\varepsilon}$ $\longrightarrow R_{\pm}$ $\text{strongly in }L^{\infty}(0,T;L^{2}(\Omega)),$ (1.10)
$\rho_{\pm}^{\varepsilon}$ $\longrightarrow\rho_{\pm}$ $\text{strongly in }L^{\infty}(0,T;L^{2}(\Omega)),$ (1.11)
$\alpha_{\pm}^{\varepsilon}$ $\longrightarrow\alpha_{\pm}$ $\text{weak-$*$ in }L^{\infty}(0,T;L^{2}(\Omega)),$ (1.12)
$u^{\varepsilon}$ $\longrightarrow u$ $\text{strongly in }L^{2}(0,T;L^{2}(\Omega)),$ (1.13)
$\nabla u^{\varepsilon}$ $\rightharpoonup\nabla u$ $\text{weakly in }L^{2}(0,T;L^{2}(\Omega)),$ (1.14)
$\operatorname{div}u^{\varepsilon}$ $\rightharpoonup 0$ $\text{weakly in }L^{2}(0,T;L^{2}(\Omega)).$ (1.15)
Remark 1.7.
Assumptions (1.9) corresponds to the standard notion of well-prepared initial data for the singular limit, as in [36]. They ensure that the initial deviations from the incompressible limit are sufficiently small to control the energy estimates uniformly in $\varepsilon$.
Remark 1.8.
The assumption (1.6) is preserved along the time by conservation of mass:
$\int_{\Omega}R_{\pm}^{\varepsilon}(t,x)\,dx=\int_{\Omega}R_{\pm,0}^{\varepsilon}(x)\,dx,\qquad\int_{\Omega}R_{\pm}(t,x)\,dx=\int_{\Omega}R_{\pm,0}(x)\,dx,\qquad\text{for all }t\in(0,T).$
The analysis relies on the relative entropy method, uniform Sobolev estimates and compactness arguments, which allow us to rigorously justify the convergence of solutions of the compressible system toward the limit system as $\varepsilon\to 0$. Our work provides a mathematical foundation for the incompressible regime of two-phase models with algebraic closure and single velocity, and clarifies the structure of the low Mach number limit in this context.
This paper is organized as follows. In Section 2, we introduce our relative entropy functional which consists of a classical contribution and an additional component based on a logarithmic comparison of the partial densities. In Section 3, we exploit the special structure of the bi-fluid system to derive several key inequalities, including an original comparison between the $L^{1}$ and $L^{2}$ norms of the partial densities and a variant of the Csiszár–Kullback–Pinsker inequality. Next, in Section 4, we prove the strong convergence of the partial densities, and finally, Section 5 completes the proof of our main theorem.
2 Relative entropy
In this section, we introduce the relative entropy functional that will be our main tool to prove Theorem 1.6.
2.1 Definition
We derive a relative entropy between a weak solution $(\rho^{\varepsilon}_{+},\rho^{\varepsilon}_{-},\alpha^{\varepsilon}_{+},u^{\varepsilon})$ of System (LABEL:system_eps), in the sense of Definition 2.1 below, and any other state $(\rho_{+},\rho_{-},\alpha_{+},u)$ of the fluid. Let us consider the relative energy functional, denoted by ${\mathcal{E}}_{1}(\rho^{\varepsilon}_{+},\rho^{\varepsilon}_{-},\alpha^{\varepsilon}_{+},u^{\varepsilon}|\rho_{+},\rho_{-},\alpha_{+},u)$ and defined as:
$\begin{split}{\mathcal{E}}_{1}\Big(\rho^{\varepsilon}_{+},\rho^{\varepsilon}_{-},\alpha^{\varepsilon}_{+},u^{\varepsilon}|\rho_{+},\rho_{-},\alpha_{+},u\Big)=&\int_{\Omega}\frac{1}{2}(\alpha^{\varepsilon}_{+}\rho^{\varepsilon}_{+}+\alpha^{\varepsilon}_{-}\rho^{\varepsilon}_{-})|u^{\varepsilon}-u|^{2}\\ &+\int_{\Omega}\frac{\alpha_{+}^{\varepsilon}}{\varepsilon^{2}}H(\rho_{+}^{\varepsilon}|\rho_{+})+\int_{\Omega}\frac{\alpha_{-}^{\varepsilon}}{\varepsilon^{2}}H(\rho_{-}^{\varepsilon}|\rho_{-}),\end{split}$ (2.1)
with, for $i=+,$-,
$H(\rho_{i}^{\varepsilon}|\rho_{i}):=\frac{1}{\gamma_{i}-1}\left((\rho_{i}^{\varepsilon})^{\gamma_{i}}-\rho_{i}^{\gamma_{i}}-\gamma_{i}\rho_{i}^{\gamma_{i}-1}(\rho_{i}^{\varepsilon}-\rho_{i})\right).$ (2.2)
In order to control the $L^{1}$ norm of the partial densities and close the Gronwall argument required for the proof of Theorem 1.6, we introduce the following functional
[…]
and we set $\ln_{+}s=\max(0,\ln s)$.
Finally, we define the total relative entropy functional:
${\mathcal{E}}\Big(\rho^{\varepsilon}_{+},\rho^{\varepsilon}_{-},\alpha^{\varepsilon}_{+},u^{\varepsilon}|\rho_{+},\rho_{-},\alpha_{+},u\Big)=({\mathcal{E}}_{1}+{\mathcal{E}}_{2})\Big(\rho^{\varepsilon}_{+},\rho^{\varepsilon}_{-},\alpha^{\varepsilon}_{+},u^{\varepsilon}|\rho_{+},\rho_{-},\alpha_{+},u\Big).$
In this definition, we always set $\alpha_{-}^{\varepsilon}=1-\alpha_{+}^{\varepsilon}$ and $\alpha_{-}=1-\alpha_{+}$.
The goal of the next subsection is to study the evolution of this entropy for the weak solution of System (LABEL:system_eps) in the following sense.
Definition 2.1.
Let us consider some initial data $(\rho_{+,0}^{\varepsilon},R_{+,0}^{\varepsilon},m_{0}^{\varepsilon})\in L^{\gamma_{+}}\times L^{\gamma_{+}}\times L^{1}$ for $\gamma_{\pm}\geqslant 2$ satisfying the compatibility conditions (1.4), where we defined $\alpha_{\pm,0}^{\varepsilon}\in[0,1]$, $\rho_{-,0}^{\varepsilon}\geqslant 0$, $\rho^{\varepsilon}_{0}\geqslant 0$ and $u_{0}^{\varepsilon}$ from $(\rho_{+,0}^{\varepsilon},R_{+,0}^{\varepsilon},m_{0}^{\varepsilon})$ (see Remark 1.2). Let $T>0$. We say that $(\rho_{\pm}^{\varepsilon},\alpha_{\pm}^{\varepsilon},u^{\varepsilon})$ is a finite energy weak solution of the compressible two-phase Navier-Stokes system (LABEL:system_eps) with algebraic closure on $(0,T)$ associated with the initial data $(\rho_{\pm,0}^{\varepsilon},\alpha_{\pm,0}^{\varepsilon},u_{0}^{\varepsilon})$ if we have
  • the regularity:
    $0\leqslant\alpha_{+}^{\varepsilon}\leqslant 1,\quad\alpha_{+}^{\varepsilon}+\alpha_{-}^{\varepsilon}=1,\quad\rho_{\pm}^{\varepsilon}\geqslant 0,\quad\rho_{+}^{\gamma_{+}}=\rho_{-}^{\gamma_{-}}\text{ a.e. in }[0,T]\times\Omega,$
    $\rho_{\pm}^{\varepsilon},\ \alpha_{\pm}^{\varepsilon}\rho_{\pm}^{\varepsilon}\in C_{\text{weak}}([0,T];L^{\gamma_{\pm}}(\Omega))\cap L^{\infty}(0,T;L^{\gamma_{\pm}}(\Omega)),$
    $\sqrt{\rho^{\varepsilon}}u^{\varepsilon}\in L^{\infty}(0,T;L^{2}(\Omega)),\quad\rho^{\varepsilon}u^{\varepsilon}\in C_{\text{weak}}([0,T];L^{\frac{2\gamma_{+}}{\gamma_{+}+1}}+L^{\frac{2\gamma_{-}}{\gamma_{-}+1}}(\Omega)),$
    $u^{\varepsilon}\in L^{2}(0,T;H^{1}_{0}(\Omega)),$
  • the continuity equation:
    $\int_{\Omega}(\alpha_{i}^{\varepsilon}\rho_{i}^{\varepsilon}\phi)(t,x)\,\mathrm{d}x-\int_{\Omega}(\alpha_{i,0}^{\varepsilon}\rho_{i,0}^{\varepsilon})(x)\phi(0,x)\,\mathrm{d}x=\int_{0}^{t}\int_{\Omega}\Big(\alpha_{i}^{\varepsilon}\rho_{i}^{\varepsilon}(\partial_{t}\phi+u^{\varepsilon}\cdot\nabla\phi)\Big)(s,x)\,\mathrm{d}x\,\mathrm{d}s,$
    for a.e. $t\in(0,T)$, for any $\phi\in\mathcal{C}^{\infty}([0,T]\times\overline{\Omega})$ and $i=+,$-,
  • the momentum equation:
    $\int_{\Omega}(\rho^{\varepsilon}u^{\varepsilon})(t,x)\cdot\phi(t,x)dx$ $-\int_{\Omega}(\rho^{\varepsilon}_{0}u^{\varepsilon}_{0}(x)\cdot\phi(0,x)dx$
    $=$ $\int_{0}^{t}\int_{\Omega}\Big(\rho^{\varepsilon}u^{\varepsilon}\cdot\partial_{t}\phi+\rho^{\varepsilon}(u^{\varepsilon}\otimes u^{\varepsilon}):\nabla\phi+\frac{1}{\varepsilon^{2}}(\rho_{+}^{\varepsilon})^{\gamma_{+}}\operatorname{div}\phi\Big)(s,x)dxds$
    $-\int_{0}^{t}\int_{\Omega}\left(\mu\nabla u^{\varepsilon}:\nabla\phi+(\mu+\lambda)\operatorname{div}u^{\varepsilon}\operatorname{div}\phi\right)(s,x)dxds$
    for a.e. $t\in(0,T)$, for any $\phi\in\mathcal{C}_{c}^{\infty}([0,T]\times\Omega;{\mathbb{R}}^{3})$,
  • the energy inequality:
    $E(t)+\mu\int_{0}^{t}\int_{\Omega}\left(\nabla u^{\varepsilon}\right)^{2}\,\mathrm{d}x\,\mathrm{d}s+\left(\lambda+\mu\right)\int_{0}^{t}\int_{\Omega}(\operatorname{div}u^{\varepsilon})^{2}\,\mathrm{d}x\,\mathrm{d}s\leqslant E(0)$ (2.4)
    for a.e. $t\in(0,T)$, where
    $E(t)=\int_{\Omega}\frac{1}{2}\Big(\rho^{\varepsilon}|u^{\varepsilon}|^{2}\Big)(t,x)\,\mathrm{d}x+\int_{\Omega}\frac{\alpha_{+}^{\varepsilon}\rho_{+}^{\varepsilon\gamma_{+}}(t,x)}{\varepsilon^{2}(\gamma_{+}-1)}\,\mathrm{d}x+\int_{\Omega}\frac{\alpha_{-}^{\varepsilon}\rho_{-}^{\varepsilon\gamma_{-}}(t,x)}{\varepsilon^{2}(\gamma_{-}-1)}\,\mathrm{d}x$ (2.5)
Remark 2.2.
The functional appearing in the energy inequality is the natural energy associated with the system. A detailed formal derivation of the corresponding energy identity for smooth solutions can be found in [29].
2.2 First part of the entropy
We begin by estimating the time derivative of ${\mathcal{E}}_{1}$ for sufficiently smooth functions $(\rho_{+},\rho_{-},\alpha_{+},u)$ without assuming that they solve System (LABEL:system_eps).
Proposition 2.3.
If $(\rho^{\varepsilon}_{+},\rho^{\varepsilon}_{-},\alpha^{\varepsilon}_{+},u^{\varepsilon})$ is a weak solution of System (LABEL:system_eps), defined as in Definition (2.1), then for all $(\rho_{+},\rho_{-},\alpha_{+},u)$ functions verifying the regularity properties of the limit solution, see Theorem 1.3, such that $p_{i}(\rho_{i})=\rho_{i}^{\gamma_{i}}$, for $i=+,$-, and $p_{+}(\rho_{+})=p_{-}(\rho_{-})$, we have
${\mathcal{E}}_{1}(t)-{\mathcal{E}}_{1}(0)\leqslant$ $-\mu\int_{0}^{t}\int_{\Omega}\left(\nabla(u^{\varepsilon}-u)\right)^{2}-(\lambda+\mu)\int_{0}^{t}\int_{\Omega}\left(\operatorname{div}(u^{\varepsilon}-u)\right)^{2}$
$+\int_{0}^{t}\int_{\Omega}\left(\alpha_{+}^{\varepsilon}\rho_{+}^{\varepsilon}+\alpha_{-}^{\varepsilon}\rho_{-}^{\varepsilon}\right)(\partial_{t}u+u\cdot\nabla u)\cdot(u-u^{\varepsilon})$
$+\int_{0}^{t}\int_{\Omega}\left(\alpha_{+}^{\varepsilon}\rho_{+}^{\varepsilon}+\alpha_{-}^{\varepsilon}\rho_{-}^{\varepsilon}\right)(u^{\varepsilon}-u)\cdot\nabla u\cdot(u-u^{\varepsilon})$
$+\frac{1}{\varepsilon^{2}}\int_{0}^{t}\int_{\Omega}\Big(\alpha_{+}^{\varepsilon}p_{+}(\rho_{+})+\alpha_{-}^{\varepsilon}p_{-}(\rho_{-})-\left(\alpha_{+}^{\varepsilon}p_{+}\left(\rho_{+}^{\varepsilon}\right)+\alpha_{-}^{\varepsilon}p_{-}\left(\rho_{-}^{\varepsilon}\right)\right)\Big)\operatorname{div}u$
$+\frac{1}{\varepsilon^{2}}\sum_{i=+,-}\int_{0}^{t}\int_{\Omega}\gamma_{i}\alpha_{i}^{\varepsilon}\rho_{i}^{\gamma_{i}-2}(\rho_{i}-\rho_{i}^{\varepsilon})\partial_{t}\rho_{i}$
$+\frac{1}{\varepsilon^{2}}\sum_{i=+,-}\int_{\Omega}\gamma_{i}\rho_{i}^{\gamma_{i}-2}(\alpha_{i}\rho_{i}u-\alpha_{i}^{\varepsilon}\rho_{i}^{\varepsilon}u^{\varepsilon})\cdot\nabla\rho_{i}$
$-\mu\int_{0}^{t}\int_{\Omega}\nabla u:\nabla(u^{\varepsilon}-u)-(\lambda+\mu)\int_{0}^{t}\int_{\Omega}\operatorname{div}u\operatorname{div}(u^{\varepsilon}-u).$
Proof.
We have
$\begin{split}{\mathcal{E}}_{1}(t)-{\mathcal{E}}_{1}(0)=&E(t)-E(0)\\ &+\frac{1}{2}\int_{\Omega}\Big(\rho^{\varepsilon}|u|^{2}\Big)(t,\cdot)-\frac{1}{2}\int_{\Omega}\Big(\rho^{\varepsilon}|u|^{2}\Big)(0,\cdot) \\ &-\int_{\Omega}\Big(\rho^{\varepsilon}u^{\varepsilon}\cdot u\Big)(t,\cdot)+\int_{\Omega}\Big(\rho^{\varepsilon}u^{\varepsilon}\cdot u\Big)(0,\cdot)\\ &-\sum_{i=+,-}\int_{\Omega}\frac{\alpha_{i}^{\varepsilon}}{\varepsilon^{2}(\gamma_{i}-1)}\Big(\rho_{i}^{\gamma_{i}}+\gamma_{i}\rho_{i}^{\gamma_{i}-1}(\rho_{i}^{\varepsilon}-\rho_{i})\Big)(t,\cdot)\\ &+\sum_{i=+,-}\int_{\Omega}\frac{\alpha_{i}^{\varepsilon}}{\varepsilon^{2}(\gamma_{i}-1)}\Big(\rho_{i}^{\gamma_{i}}+\gamma_{i}\rho_{i}^{\gamma_{i}-1}(\rho_{i}^{\varepsilon}-\rho_{i})\Big)(0,\cdot).\end{split}$ (2.7)
Let us consider mass equations with $\phi=\frac{|u|^{2}}{2}$ and add the two, we get:
$\frac{1}{2}\int_{\Omega}\Big(\rho^{\varepsilon}|u|^{2}\Big)(t,\cdot)-\frac{1}{2}\int_{\Omega}\Big(\rho^{\varepsilon}|u|^{2}\Big)(0,\cdot)$  $=\frac{1}{2}\int_{0}^{t}\!\!\!\int_{\Omega}\rho^{\varepsilon}\left(\partial_{t}|u|^{2}+u^{\varepsilon}\cdot\nabla|u|^{2}\right)$
$=\int_{0}^{t}\!\!\!\int_{\Omega}\rho^{\varepsilon}u\cdot\Big(\partial_{t}u+(u^{\varepsilon}\cdot\nabla)u\Big).$
Let us notice that we need, here and below, to check that all terms are well defined by using the regularity properties of the functions depending on $\varepsilon$ (see Definition 2.1) and the smooth functions (see Theorem 1.3).
Let us now use the momentum equation by taking $\phi=u$, we get:
$\int_{\Omega}\Big(\rho^{\varepsilon}u^{\varepsilon}\cdot u\Big)(t,\cdot)+\int_{\Omega}\Big(\rho^{\varepsilon}u^{\varepsilon}\cdot u\Big)(0,\cdot)=-\int_{0}^{t}\!\!\!\int_{\Omega}\rho^{\varepsilon}u^{\varepsilon}\cdot(\partial_{t}u+(u^{\varepsilon}\cdot\nabla)u)\\ -\frac{1}{\varepsilon^{2}}\int_{0}^{t}\!\!\!\int_{\Omega}\rho_{+}^{\varepsilon\gamma_{+}}\operatorname{div}u+\mu\int_{0}^{t}\!\!\!\int_{\Omega}\nabla u^{\varepsilon}:\nabla u+(\mu+\lambda)\int_{0}^{t}\!\!\!\int_{\Omega}\left(\operatorname{div}u^{\varepsilon}\right)(\operatorname{div}u),$
where we can replace
$-\frac{1}{\varepsilon^{2}}\int_{0}^{t}\!\!\!\int_{\Omega}\rho_{+}^{\varepsilon\gamma_{+}}\operatorname{div}u$ $=-\frac{1}{\varepsilon^{2}}\int_{0}^{t}\!\!\!\int_{\Omega}(\alpha_{+}^{\varepsilon}+\alpha_{-}^{\varepsilon})\rho_{+}^{\varepsilon\gamma_{+}}\operatorname{div}u$
$=-\frac{1}{\varepsilon^{2}}\int_{0}^{t}\!\!\!\int_{\Omega}(\alpha_{+}^{\varepsilon}p_{+}(\rho_{+}^{\varepsilon})+\alpha_{-}^{\varepsilon}p_{-}(\rho_{-}^{\varepsilon}))\operatorname{div}u.$
We can also rewrite
$\int_{\Omega}\rho^{\varepsilon}(\partial_{t}u+(u^{\varepsilon}\cdot\nabla)u)\cdot(u-u^{\varepsilon})=$ $\int_{\Omega}\rho^{\varepsilon}(\partial_{t}u+(u\cdot\nabla)u)\cdot(u-u^{\varepsilon})$
$+\rho^{\varepsilon}\Big(((u^{\varepsilon}-u)\cdot\nabla)u\Big)\cdot(u-u^{\varepsilon}).$
By the equality of the pressures and the fact that $\alpha_{+}^{\varepsilon}+\alpha_{-}^{\varepsilon}=1$, we have
$-\sum_{i=+,-}\int_{\Omega}\frac{\alpha_{i}^{\varepsilon}}{\varepsilon^{2}(\gamma_{i}-1)}\Big(\rho_{i}^{\gamma_{i}}-\gamma_{i}\rho_{i}^{\gamma_{i}-1}\rho_{i}\Big)=\sum_{i=+,-}\int_{\Omega}\frac{\alpha_{i}^{\varepsilon}}{\varepsilon^{2}}\rho_{i}^{\gamma_{i}}=\int_{\Omega}\frac{1}{\varepsilon^{2}}\rho_{+}^{\gamma_{+}}$
hence
$\sum_{i=+,-}\int_{\Omega}\frac{\alpha_{i}^{\varepsilon}}{\varepsilon^{2}(\gamma_{i}-1)}\Big(\rho_{i}^{\gamma_{i}}-\gamma_{i}\rho_{i}^{\gamma_{i}-1}\rho_{i}\Big)(t,\cdot)+\int_{\Omega}\frac{\alpha_{i}^{\varepsilon}}{\varepsilon^{2}(\gamma_{i}-1)}\Big(\rho_{i}^{\gamma_{i}}-\gamma_{i}\rho_{i}^{\gamma_{i}-1}\rho_{i})\Big)(0,\cdot)\\ =\int_{0}^{t}\!\!\!\int_{\Omega}\frac{1}{\varepsilon^{2}}(\alpha_{+}^{\varepsilon}+\alpha_{-}^{\varepsilon})\partial_{t}\rho_{+}^{\gamma_{+}}=\sum_{i=+,-}\int_{0}^{t}\!\!\!\int_{\Omega}\frac{1}{\varepsilon^{2}}\alpha_{i}^{\varepsilon}\gamma_{i}\rho_{i}^{\gamma-1}\partial_{t}\rho_{i}$
For $i=+,$-, remark now that the mass equation with $\phi=\rho_{i}^{\gamma_{i}-1}$ gives
$-\int_{\Omega}\frac{\gamma_{i}\alpha_{i}^{\varepsilon}}{\varepsilon^{2}(\gamma_{i}-1)}\rho_{i}^{\gamma_{i}-1}\rho_{i}^{\varepsilon}(t,\cdot)+\int_{\Omega}\frac{\gamma_{i}\alpha_{i}^{\varepsilon}}{\varepsilon^{2}(\gamma_{i}-1)}\rho_{i}^{\gamma_{i}-1}\rho_{i}^{\varepsilon}(0,\cdot)$ $=-\int_{0}^{t}\!\!\!\int_{\Omega}\frac{\gamma_{i}\alpha_{i}^{\varepsilon}\rho_{i}^{\varepsilon}\rho_{i}^{\gamma_{i}-2}}{\varepsilon^{2}}\Big(\partial_{t}\rho_{i}+(u^{\varepsilon}\cdot\nabla)\rho_{i}\Big).$
We then have obtained
$-\sum_{i=+,-}\int_{\Omega}\frac{\alpha_{i}^{\varepsilon}}{\varepsilon^{2}(\gamma_{i}-1)}\Big(\rho_{i}^{\gamma_{i}}+\gamma_{i}\rho_{i}^{\gamma_{i}-1}(\rho_{i}^{\varepsilon}-\rho_{i})\Big)(t,\cdot)$
$+\sum_{i=+,-}\int_{\Omega}\frac{\alpha_{i}^{\varepsilon}}{\varepsilon^{2}(\gamma_{i}-1)}\Big(\rho_{i}^{\gamma_{i}}+\gamma_{i}\rho_{i}^{\gamma_{i}-1}(\rho_{i}^{\varepsilon}-\rho_{i})\Big)(0,\cdot)$ $=\frac{1}{\varepsilon^{2}}\sum_{i=+,-}\int_{0}^{t}\!\!\!\int_{\Omega}\gamma_{i}\alpha_{i}^{\varepsilon}\rho_{i}^{\gamma_{i}-2}(\rho_{i}-\rho_{i}^{\varepsilon})\partial_{t}\rho_{i}$
$+\frac{1}{\varepsilon^{2}}\sum_{i=+,-}\int_{0}^{t}\!\!\!\int_{\Omega}\gamma_{i}\rho_{i}^{\gamma_{i}-2}(\alpha_{i}\rho_{i}u-\alpha_{i}^{\varepsilon}\rho_{i}^{\varepsilon}u^{\varepsilon})\cdot\nabla\rho_{i}$
$-\frac{1}{\varepsilon^{2}}\sum_{i=+,-}\int_{0}^{t}\!\!\!\int_{\Omega}\gamma_{i}\rho_{i}^{\gamma_{i}-2}\alpha_{i}\rho_{i}u\cdot\nabla\rho_{i}$
Using again that $p_{+}(\rho_{+})=p_{-}(\rho_{-})$ and $\alpha_{+}+\alpha_{-}=1$, we can rewrite the last term
$-\frac{1}{\varepsilon^{2}}\sum_{i=+,-}\int_{0}^{t}\!\!\!\int_{\Omega}\gamma_{i}\rho_{i}^{\gamma_{i}-2}\alpha_{i}\rho_{i}u\cdot\nabla\rho_{i}=$ $-\frac{1}{\varepsilon^{2}}\sum_{i=+,-}\int_{0}^{t}\!\!\!\int_{\Omega}\alpha_{i}u\cdot\nabla p_{i}(\rho_{i})=-\frac{1}{\varepsilon^{2}}\int_{0}^{t}\!\!\!\int_{\Omega}u\cdot\nabla p_{+}(\rho_{+})$
$=$ $\frac{1}{\varepsilon^{2}}\int_{0}^{t}\!\!\!\int_{\Omega}p_{+}(\rho_{+})\operatorname{div}u$
$=$ $\frac{1}{\varepsilon^{2}}\int_{0}^{t}\!\!\!\int_{\Omega}(\alpha_{+}^{\varepsilon}p_{+}(\rho_{+})+\alpha_{-}^{\varepsilon}p_{-}(\rho_{-}))\operatorname{div}u$
where we have used that $u$ satisfies the Dirichlet boundary condition.
Therefore, using the energy estimate (2.4) satisfied by $(\rho_{+}^{\varepsilon},\rho_{-}^{\varepsilon},\alpha_{+}^{\varepsilon},u^{\varepsilon})$ and plugging all the previous computations into (2.7), we obtain:
[…]
Finally, we end the proof by noticing that by virtue of monotonicity of the operators $\nabla$ and $\operatorname{div}$, namely, we write
$-\nabla u^{\varepsilon}:\nabla(u^{\varepsilon}-u)$ $=-\left(\nabla(u^{\varepsilon}-u)\right)^{2}-\nabla u:\nabla(u^{\varepsilon}-u),$
and
$-\operatorname{div}u^{\varepsilon}\operatorname{div}(u^{\varepsilon}-u)$ $=-\left(\operatorname{div}(u^{\varepsilon}-u)\right)^{2}-\operatorname{div}u\operatorname{div}(u^{\varepsilon}-u).$
Let us define the remainder $\mathcal{R}$} as almost all terms on the right-hand side of Proposition 2.3:
${\mathcal{R}}\left(\rho^{\varepsilon}_{+},\rho^{\varepsilon}_{-},\alpha^{\varepsilon}_{+},u^{\varepsilon}|\rho_{+},\rho_{-},\alpha_{+},u\right):=\int_{0}^{t}\int_{\Omega}\left(\alpha_{+}^{\varepsilon}\rho_{+}^{\varepsilon}+\alpha_{-}^{\varepsilon}\rho_{-}^{\varepsilon}\right)(\partial_{t}u+u\cdot\nabla u)\cdot(u-u^{\varepsilon})$ (2.8a)
$+\int_{0}^{t}\int_{\Omega}\left(\alpha_{+}^{\varepsilon}\rho_{+}^{\varepsilon}+\alpha_{-}^{\varepsilon}\rho_{-}^{\varepsilon}\right)(u^{\varepsilon}-u)\cdot\nabla u\cdot(u-u^{\varepsilon})$ (2.8b)
$+\frac{1}{\varepsilon^{2}}\int_{0}^{t}\int_{\Omega}\Big(\alpha_{+}^{\varepsilon}p_{+}(\rho_{+})+\alpha_{-}^{\varepsilon}p_{-}(\rho_{-})-\left(\alpha_{+}^{\varepsilon}p_{+}\left(\rho_{+}^{\varepsilon}\right)+\alpha_{-}^{\varepsilon}p_{-}\left(\rho_{-}^{\varepsilon}\right)\right)\Big)\operatorname{div}u$ (2.8c)
$+\frac{1}{\varepsilon^{2}}\sum_{i=+,-}\int_{0}^{t}\int_{\Omega}\gamma_{i}\alpha_{i}^{\varepsilon}\rho_{i}^{\gamma_{i}-2}(\rho_{i}-\rho_{i}^{\varepsilon})\partial_{t}\rho_{i}$ (2.8d)
$+\frac{1}{\varepsilon^{2}}\sum_{i=+,-}\int_{\Omega}\gamma_{i}\rho_{i}^{\gamma_{i}-2}(\alpha_{i}\rho_{i}u-\alpha_{i}^{\varepsilon}\rho_{i}^{\varepsilon}u^{\varepsilon})\cdot\nabla\rho_{i}$ (2.8e)
$-\mu\int_{0}^{t}\int_{\Omega}\nabla u:\nabla(u^{\varepsilon}-u)-(\lambda+\mu)\int_{0}^{t}\int_{\Omega}\operatorname{div}u\operatorname{div}(u^{\varepsilon}-u).$ (2.8f)
We now consider $(\rho_{+},\rho_{-},\alpha_{+},u)$ as a solution of System (1.5), verifying the regularity properties of the limit solution, see Theorem 1.3. We can reformulate Proposition 2.3.
Proposition 2.4.
If $(\rho^{\varepsilon}_{+},\rho^{\varepsilon}_{-},\alpha^{\varepsilon}_{+},u^{\varepsilon})$ is a weak solution of System (LABEL:system_eps), defined as in Definition 2.1, and if $(\rho_{+},\rho_{-},\alpha_{+},u)$ is a strong solution of System (1.5), then
${\mathcal{R}}\left(\rho^{\varepsilon}_{+},\rho^{\varepsilon}_{-},\alpha^{\varepsilon}_{+},u^{\varepsilon}|\rho_{+},\rho_{-},\alpha_{+},u\right)=$ $\mu\int_{0}^{t}\int_{\Omega}\frac{\rho^{\varepsilon}-\rho}{\rho}\Delta u(u-u^{\varepsilon})$
$+\int_{0}^{t}\int_{\Omega}\rho^{\varepsilon}(u-u^{\varepsilon})\cdot\nabla u\cdot(u^{\varepsilon}-u)$
$-\int_{0}^{t}\int_{\Omega}\frac{\rho^{\varepsilon}}{\rho}\nabla\Pi(u-u^{\varepsilon}).$
Proof.
First, let us note that because we consider a solution of System (1.5), we have $\operatorname{div}u=0$, $\partial_{t}\rho_{i}=0$ and $\nabla\rho_{i}=0$, therefore (2.8c), (2.8d) and (2.8e) in the remainder $\mathcal{R}$} are cancelled and we only have
${\mathcal{R}}\left(\rho^{\varepsilon}_{+},\rho^{\varepsilon}_{-},\alpha^{\varepsilon}_{+},u^{\varepsilon}|\rho_{+},\rho_{-},\alpha_{+},u\right)=$ $\int_{0}^{t}\int_{\Omega}\left(\alpha_{+}^{\varepsilon}\rho_{+}^{\varepsilon}+\alpha_{-}^{\varepsilon}\rho_{-}^{\varepsilon}\right)(\partial_{t}u+u\cdot\nabla u)\cdot(u-u^{\varepsilon})$
$+\int_{0}^{t}\int_{\Omega}\left(\alpha_{+}^{\varepsilon}\rho_{+}^{\varepsilon}+\alpha_{-}^{\varepsilon}\rho_{-}^{\varepsilon}\right)(u^{\varepsilon}-u)\cdot\nabla u\cdot(u-u^{\varepsilon})$
$+\int_{0}^{t}\mu\int_{\Omega}\nabla u:\nabla(u-u^{\varepsilon}).$
Denoting $\rho=\alpha_{+}\rho_{+}+\alpha_{-}\rho_{-}$ and $\rho_{\varepsilon}=\alpha_{+}^{\varepsilon}\rho_{+}^{\varepsilon}+\alpha_{-}^{\varepsilon}\rho_{-}^{\varepsilon}$, we use the momentum equation verified by $(\rho_{+},\rho_{-},\alpha_{+},u)$ to get:
${\mathcal{R}}\left(\rho^{\varepsilon}_{+},\rho^{\varepsilon}_{-},\alpha^{\varepsilon}_{+},u^{\varepsilon}|\rho_{+},\rho_{-},\alpha_{+},u\right)=\int_{0}^{t}\int_{\Omega}\frac{\rho^{\varepsilon}}{\rho}\mu\Delta u(u-u^{\varepsilon})+\mu\int_{0}^{t}\int_{\Omega}\nabla u:\nabla(u-u^{\varepsilon})\\ -\int_{0}^{t}\int_{\Omega}\frac{\rho^{\varepsilon}}{\rho}\nabla\Pi\cdot(u-u^{\varepsilon})+\int_{0}^{t}\int_{\Omega}\rho^{\varepsilon}(u-u^{\varepsilon})\cdot\nabla u\cdot(u^{\varepsilon}-u).$
Finally, we end the proof by noticing that
$\int_{\Omega}\mu\frac{\rho^{\varepsilon}}{\rho}\Delta u(u-u^{\varepsilon})+\int_{\Omega}\mu\nabla u:\nabla(u-u^{\varepsilon})$ $=\mu\int_{\Omega}\frac{\rho^{\varepsilon}}{\rho}\Delta u(u-u^{\varepsilon})-\mu\int_{\Omega}\Delta u(u-u^{\varepsilon})$
$=\mu\int_{\Omega}\frac{\rho^{\varepsilon}-\rho}{\rho}\Delta u(u-u^{\varepsilon}).$
2.3 Second part of the entropy
We now study the evolution of the entropy component (2.3) for weak solutions of System (LABEL:system_eps).
Proposition 2.5.
If $(\rho^{\varepsilon}_{+},\rho^{\varepsilon}_{-},\alpha^{\varepsilon}_{+},u^{\varepsilon})$ is a weak solution of System (LABEL:system_eps), defined as in Definition (2.1), and if $(\rho_{+},\rho_{-},\alpha_{+},u)$ is a solution of System (1.5), verifying the regularity properties of the limit solution, see Theorem 1.3, then there exits $C>0$ depending on $\|\nabla R_{i}\|_{L^{3}}$ such that
$\begin{split}\int_{\Omega}R_{i}^{\varepsilon}\ln\Big(\frac{R_{i}^{\varepsilon}}{R_{i}}\Big)(t)-\int_{\Omega}R_{i}^{\varepsilon}\ln\Big(\frac{R_{i}^{\varepsilon}}{R_{i}}\Big)(0)&\leqslant C\sum_{i=+,-}\int_{0}^{t}\|R_{i}^{\varepsilon}-R_{i}\|_{L^{2}(\Omega)}\|\nabla(u^{\varepsilon}-u)\|_{L^{2}(\Omega)}\\ &\quad+C\sum_{i=+,-}\int_{0}^{t}\|R_{i}^{\varepsilon}-R_{i}\|_{L^{2}(\Omega)}\|\operatorname{div}(u-u^{\varepsilon})\|_{L^{2}(\Omega)}.\end{split}$ (2.10)
Proof.
We want to take $\phi=\ln R_{\pm}^{\varepsilon}$ in the continuity equation of Definition 2.1. To justify the use of a possibly non-admissible test function, we appeal to the theory of renormalized solutions. Since
$u^{\varepsilon}\in L^{2}(0,T;H^{1}(\Omega)),$
we regularize the continuity equation by convolution. At the regularized level, the quantities are smooth enough to perform the computations rigorously. This procedures is classical, in particular, it is used by P.-L. Lions in the barotropic compressible Navier–Stokes system to obtain compactness of the density. It requires
$\rho\in L^{2}(0,T;L^{2}(\Omega)),$
if one wishes to avoid the truncation–regularization method of E. Feireisl.
We have
$\int_{\Omega}(R_{i}^{\varepsilon}\ln R_{i}^{\varepsilon})(t,x)\,\mathrm{d}x-\int_{\Omega}R_{i,0}^{\varepsilon}(x)\ln R_{i,0}^{\varepsilon}(x)\,\mathrm{d}x$ $=\int_{0}^{t}\int_{\Omega}\Big(R_{i}^{\varepsilon}(\partial_{t}\ln R_{i}^{\varepsilon}+u^{\varepsilon}\cdot\nabla\ln R_{i}^{\varepsilon})\Big)(s,x)\,\mathrm{d}x\,\mathrm{d}s,$
for a.a. $t\in(0,T)$ and $i=+,$-. Therefore
$\int_{\Omega}(R_{i}^{\varepsilon}\ln R_{i}^{\varepsilon})(t,x)\,\mathrm{d}x$ $-\int_{\Omega}R_{i,0}^{\varepsilon}(x)\ln R_{i,0}^{\varepsilon}(x)\,\mathrm{d}x-\Big(\int_{\Omega}R_{i}^{\varepsilon}(t,x)\,\mathrm{d}x-\int_{\Omega}R_{i,0}^{\varepsilon}(x)\,\mathrm{d}x\Big)$
$+\int_{0}^{t}\int_{\Omega}R_{i}^{\varepsilon}\operatorname{div}u^{\varepsilon}(s,x)\,\mathrm{d}x\,\mathrm{d}s=0.$ (2.11)
We multiply now the mass equation of $R_{\pm}$ by $R_{\pm}^{\varepsilon}/R_{\pm}$ and we can write, for $i=+,$-,
$\int_{\Omega}R_{i}^{\varepsilon}(t,x)\,\mathrm{d}x-\int_{\Omega}R_{i,0}^{\varepsilon}(x)\,\mathrm{d}x$ $=\int_{0}^{t}\int_{\Omega}\frac{R_{i}^{\varepsilon}}{R_{i}}\Big(\partial_{t}R_{i}+u\cdot\nabla R_{i})\Big)(s,x)\,\mathrm{d}x\,\mathrm{d}s.$
Therefore
$\begin{split}\int_{\Omega}R_{i}^{\varepsilon}(t,x)\,\mathrm{d}x-\int_{\Omega}R_{i,0}^{\varepsilon}(x)\,\mathrm{d}x&=\int_{0}^{t}\int_{\Omega}R_{i}^{\varepsilon}(\partial_{t}\ln R_{i}+u^{\varepsilon}\cdot\nabla\ln_{+}R_{i})(s,x)\,\mathrm{d}x\,\mathrm{d}s\\ &\quad-\int_{0}^{t}\int_{\Omega}\Big(R_{i}^{\varepsilon}(u^{\varepsilon}-u)\nabla\ln R_{i}\Big)(s,x)\,\mathrm{d}x\,\mathrm{d}s\\ &\quad+\int_{0}^{t}\int_{\Omega}\Big(R_{i}^{\varepsilon}\operatorname{div}u\Big)(s,x)\,\mathrm{d}x\,\mathrm{d}s.\end{split}$
Then we use the mass equation of $R_{i}^{\varepsilon}$ in the weak sense taking $\phi=\ln R_{i}$, we obtain
$\begin{split}\int_{\Omega}R_{i}^{\varepsilon}(t,x)\,\mathrm{d}x-\int_{\Omega}R_{i,0}^{\varepsilon}(x)\,\mathrm{d}x&=\int_{\Omega}(R_{i}^{\varepsilon}\ln R_{i})(t,x)\,\mathrm{d}x-\int_{\Omega}R_{i,0}^{\varepsilon}(x)\ln R_{i,0}(x)\,\mathrm{d}x\\ &\quad-\int_{0}^{t}\int_{\Omega}\Big(R_{i}^{\varepsilon}(u^{\varepsilon}-u)\nabla\ln R_{i}\Big)(s,x)\,\mathrm{d}x\,\mathrm{d}s\\ &\quad+\int_{0}^{t}\int_{\Omega}\Big(R_{i}^{\varepsilon}\operatorname{div}u\Big)(s,x)\,\mathrm{d}x\,\mathrm{d}s.\end{split}$ (2.12)
Adding (2.11) and (2.12) we get
$\int_{\Omega}\Big(R_{i}^{\varepsilon}\ln\frac{R_{i}^{\varepsilon}}{R_{i}}\Big)(t,x)\,\mathrm{d}x-\int_{\Omega}R_{i,0}^{\varepsilon}(x)\ln\frac{R_{i,0}^{\varepsilon}(x)}{R_{i,0}(x)}\,\mathrm{d}x+\int_{0}^{t}\int_{\Omega}\Big((R_{i}^{\varepsilon}-R_{i})(u^{\varepsilon}-u)\nabla\ln R_{i}\Big)(s,x)\,\mathrm{d}x\,\mathrm{d}s$
$\qquad+\int_{0}^{t}\int_{\Omega}\Big(R_{i}(u^{\varepsilon}-u)\nabla\ln R_{i}\Big)(s,x)\,\mathrm{d}x\,\mathrm{d}s+\int_{0}^{t}\int_{\Omega}\Big(R_{i}^{\varepsilon}\operatorname{div}(u^{\varepsilon}-u)\Big)(s,x)\,\mathrm{d}x\,\mathrm{d}s=0.$
Finally, we obtain
$\int_{\Omega}\Big(R_{i}^{\varepsilon}\ln\frac{R_{i}^{\varepsilon}}{R_{i}}\Big)(t,x)\,\mathrm{d}x-\int_{\Omega}R_{i,0}^{\varepsilon}(x)\ln\frac{R_{i,0}^{\varepsilon}(x)}{R_{i,0}(x)}\,\mathrm{d}x=-\int_{0}^{t}\int_{\Omega}\Big((R_{i}^{\varepsilon}-R_{i})(u^{\varepsilon}-u)\nabla\ln R_{i}\Big)(s,x)\,\mathrm{d}x\,\mathrm{d}s$
$\qquad-\int_{0}^{t}\int_{\Omega}\Big((R_{i}^{\varepsilon}-R_{i})\operatorname{div}(u^{\varepsilon}-u)\Big)(s,x)\,\mathrm{d}x\,\mathrm{d}s.$
Using Hölder inequality for the first term we get
$\int_{\Omega}\lvert(R_{\pm}^{\varepsilon}-R_{\pm})(u^{\varepsilon}-u)\nabla\ln_{+}R_{\pm}\rvert$ $\leqslant\|R_{\pm}^{\varepsilon}-R_{\pm}\|_{L^{2}(\Omega)}\lVert\nabla\ln_{+}R_{\pm}\rVert_{L^{3}(\Omega)}\lVert u-u^{\varepsilon}\rVert_{L^{6}(\Omega)}.$
We use Poincaré inequality to obtain
$\int_{\Omega}\lvert(R_{\pm}^{\varepsilon}-R_{\pm})(u^{\varepsilon}-u)\nabla\ln_{+}R_{\pm}\rvert$ $\leqslant C\|R_{\pm}^{\varepsilon}-R_{\pm}\|_{L^{2}(\Omega)}\lVert\nabla(u-u^{\varepsilon})\rVert_{L^{2}(\Omega)}^{2}.$
And similarly we get for the second term
$\int_{\Omega}\lvert(R_{\pm}^{\varepsilon}-R_{\pm})\operatorname{div}(u^{\varepsilon}-u)\rvert$ $\leqslant\|R_{\pm}^{\varepsilon}-R_{\pm}\|_{L^{2}(\Omega)}\|\operatorname{div}(u-u^{\varepsilon})\|_{L^{2}(\Omega)}.$
and we reach the desired conclusion. ∎
Corollary 2.6.
If $(\rho^{\varepsilon}_{+},\rho^{\varepsilon}_{-},\alpha^{\varepsilon}_{+},u^{\varepsilon})$ is a weak solution of System (LABEL:system_eps), defined as in Definition (2.1), and if $(\rho_{+},\rho_{-},\alpha_{+},u)$ is a solution of System (1.5), verifying the regularity properties of the limit solution, see Theorem 1.3, then there exits $C>0$ depending on $\|\nabla R_{i}\|_{L^{3}}$ such that
${\mathcal{E}}_{2}(t)-{\mathcal{E}}_{2}(0)\leqslant C\sum_{i=+,-}\int_{0}^{t}\|R_{i}^{\varepsilon}-R_{i}\|_{L^{2}(\Omega)}\Big(\|\nabla(u^{\varepsilon}-u)\|_{L^{2}(\Omega)}+\|\operatorname{div}(u-u^{\varepsilon})\|_{L^{2}(\Omega)}\Big)+C.$
Proof.
First, note that we have
$f\ln_{+}f=f\ln f-f\ln f1_{f\in(0,1)}$
Using the fact that $x\ln x1_{(0,1)}\leqslant C$, we obtain the following inequality (see [20] for similar inequality)
$f\ln_{+}f\leqslant f\ln f+C.$
It implies that
$R_{i}^{\varepsilon}\ln_{+}\Big(\frac{R_{i}^{\varepsilon}}{R_{i}}\Big)(t)\leqslant R_{i}^{\varepsilon}\ln\Big(\frac{R_{i}^{\varepsilon}}{R_{i}}\Big)(t)+CR_{i}$
and then
$\int_{\Omega}R_{i}^{\varepsilon}\ln_{+}\Big(\frac{R_{i}^{\varepsilon}}{R_{i}}\Big)(t)$ $-\int_{\Omega}R_{i}^{\varepsilon}\ln_{+}\Big(\frac{R_{i}^{\varepsilon}}{R_{i}}\Big)(0)$
$\leqslant$ $\int_{\Omega}R_{i}^{\varepsilon}\ln\Big(\frac{R_{i}^{\varepsilon}}{R_{i}}\Big)(t)-\int_{\Omega}R_{i}^{\varepsilon}\ln\Big(\frac{R_{i}^{\varepsilon}}{R_{i}}\Big)(0)$
$+\int_{\Omega}R_{i}^{\varepsilon}\ln\Big(\frac{R_{i}^{\varepsilon}}{R_{i}}\Big)(0)-\int_{\Omega}R_{i}^{\varepsilon}\ln_{+}\Big(\frac{R_{i}^{\varepsilon}}{R_{i}}\Big)(0)+C\int R_{i}(t)$
$\leqslant C\int_{0}^{t}\|R_{i}^{\varepsilon}-R_{i}\|_{L^{2}}\Big(\|\nabla(u^{\varepsilon}-u)\|_{L^{2}(\Omega)}+\|\operatorname{div}(u-u^{\varepsilon})\|_{L^{2}(\Omega)}\Big)+C\int R_{i}(0)$
where we have used Remark 1.8 and (1.6).
3 Key estimates for the convergence analysis
In this section, we will always denote by $(\rho^{\varepsilon}_{+},\rho^{\varepsilon}_{-},\alpha^{\varepsilon}_{+},u^{\varepsilon})$ a weak solution of System (LABEL:system_eps) in the sense of Definition 2.1 and $(\rho_{+},\rho_{-},\alpha_{+},u)$ a solution of System (1.5), verifying the regularity properties of the limit solution, see Theorem 1.3.
We first recall some known inequalities about the function $H$ given by (2.2) (see [36], [17], [43]). Assume some $\gamma>1$. For any compact $K\subset(0,+\infty)$ there exists two constants $c_{1}$ and $c_{2}$ such that for any $\rho\geqslant 0$ and for any $r\in K$,
$c_{1}\left(\lvert\rho-r\rvert^{2}1_{\lvert\rho-r\rvert<1}+\lvert\rho-r\rvert^{\gamma}1_{\lvert\rho-r\rvert\geqslant 1}\right)\leqslant H(\rho|r),$
$H(\rho|r)\leqslant c_{2}\left(\lvert\rho-r\rvert^{2}1_{\lvert\rho-r\rvert<1}+\lvert\rho-r\rvert^{\gamma}1_{\lvert\rho-r\rvert\geqslant 1}\right).$
Let us define, for $i=+,$-, $B_{i}^{1}=\{\lvert\rho_{i}^{\varepsilon}-\rho_{i}\rvert\leqslant 1\}$ and $B_{i}^{2}=\{\lvert\rho_{i}^{\varepsilon}-\rho_{i}\rvert\geqslant 1\}$. Therefore, we have, for $i=+,$-
$\int_{B_{i}^{1}}\alpha_{i}^{\varepsilon}\lvert\rho_{i}^{\varepsilon}-\rho_{i}\rvert^{2}\leqslant C\int_{B_{i}^{1}}\alpha_{i}^{\varepsilon}H(\rho_{i}^{\varepsilon}\rvert\rho_{i})\leqslant C\varepsilon^{2}{\mathcal{E}}_{1},$ (3.1)
$\int_{B_{i}^{2}}\alpha_{i}^{\varepsilon}\lvert\rho_{i}^{\varepsilon}-\rho_{i}\rvert^{\gamma_{i}}\leqslant C\int_{B_{i}^{2}}\alpha_{i}^{\varepsilon}H(\rho_{i}^{\varepsilon}\rvert\rho_{i})\leqslant C\varepsilon^{2}{\mathcal{E}}_{1}.$ (3.2)
3.1 A key $L^{2}$-estimate for the partial densities
In this subsection, we will use the bi-fluid structure to obtain a control of the volume fractions by the partial densities, which will allows us to get an estimate of the $L^{2}$ norm of $R_{\pm}^{\varepsilon}-R_{\pm}$ by the $L^{1}$ norm and the entropy.
We first prove the following fundamental result:
Lemma 3.1.
Assume that there exists a constant $c>0$ such that $R_{+}^{\varepsilon}>c$. Therefore, we have
$\lvert\alpha_{+}^{\varepsilon}-\alpha_{+}\rvert\leqslant C\lvert R_{+}^{\varepsilon}-R_{+}\rvert+C\lvert R_{-}^{\varepsilon}-R_{-}\rvert$. (3.3)
By symmetry, we have the similar result for $|\alpha_{-}^{\varepsilon}-\alpha_{-}|$ when $R_{-}^{\varepsilon}>c$. To prove this result, we first establish a key lemma, which provides the necessary technical tools.
Lemma 3.2.
Let $\gamma>0$. For all $d_{1},d_{2}>0$ such that $d_{2}\in[\underline{c},\overline{c}]\subset{\mathbb{R}}^{+}_{*}$, there exists a unique solution $(a_{1},a_{2})\in(0,1)$ that satisfies
$d_{1}a_{1}^{\gamma}+(a_{1}-1)=0,$ (3.4)
$d_{2}a_{2}^{\gamma}+(a_{2}-1)=0.$ (3.5)
Moreover, there exists a constant $C$ (that only depends on $\gamma,\overline{c},\underline{c}$) such that
$\lvert a_{1}-a_{2}\rvert\leqslant C\lvert d_{1}-d_{2}\rvert$. (3.6)
We start by proving Lemma 3.2.
Proof of Lemma 3.2.
For each $i=1,2$, we define the function $F_{d_{i}}:\mathbb{R}_{+}\to\mathbb{R}$ by
$F_{d_{i}}(x):=d_{i}x^{\gamma}+x-1.$
The function $F_{d_{i}}$ is strictly increasing with respect to $x$. Moreover, we have $F_{d_{i}}(0)=-1$ and $F_{d_{i}}(1)=d_{i}>0$. By continuity, there exists a unique solution $x_{i}\in(0,1)$ to the equation $F_{d_{i}}(x)=0$. We denote this solution by $a_{i}$ for $i=1,2$.
We then define the function $g$ by setting $a_{i}=g(d_{i})$ for all $i$. The monotonicity of $F_{d_{i}}$ with respect to $x$ implies that $g$ is decreasing with respect to $d_{i}$.
First, consider the case where $d_{1}<\frac{\underline{c}}{2}$ or $d_{1}>2\overline{c}$. Since $d_{2}\in[\underline{c},\overline{c}]$, there exists a constant $C>0$ sufficiently large such that
$|d_{1}-d_{2}|\geqslant 1/C.$
Moreover, since $a_{1},a_{2}\in(0,1)$, it follows that
$|a_{1}-a_{2}|\leqslant 1\leqslant C|d_{1}-d_{2}|.$
Now consider the case where $d_{1}\in[\frac{\underline{c}}{2},2\overline{c}]$. Since the function $g$ is decreasing, it follows that $a_{1}\in[g(2\overline{c}),g(\frac{\underline{c}}{2})]$. We then introduce the compact set $K=[\frac{\underline{c}}{2},2\overline{c}]\times[g(2\overline{c}),g(\frac{\underline{c}}{2})]$ and and define the function $F:K\to\mathbb{R}$ by $F(d,x)=dx^{\gamma}+x-1$. The function $F$ is of class $C^{\infty}$ on $K$.
For any $(d_{i},x_{i})\in K$ satisfying $F(d_{i},x_{i})=0$ and $\partial_{x}F(d_{i},x_{i})\neq 0$, the Implicit Function Theorem ensures the existence of an open neighbourhood $U_{i,\mathrm{loc}}$ of $d_{i}$ and a unique function $g_{i,\mathrm{loc}}:U_{i,\mathrm{loc}}\to\mathbb{R}$ such that
$x_{i}=g_{i,\mathrm{loc}}(d_{i})\quad\text{and}\quad F\bigl(d,g_{i,\mathrm{loc}}(d)\bigr)=0\quad\text{for all }d\in U_{i,\mathrm{loc}}.$
Since the solution $(d_{i},x_{i})$ of the equation $F(d_{i},x)=0$ is unique, it follows from the uniqueness that
$g_{i,\mathrm{loc}}=g.$
From the Implicit Function Theorem, we also have
$g_{i,\mathrm{loc}}^{\prime}(d)=-\,\frac{\partial_{d}F(d,g(d))}{\partial_{x}F(d,g(d))}=-\,\frac{x^{\gamma}}{\gamma dx^{\gamma-1}+1},\qquad\text{with }x=g(d).$
Since $(d,x)\in K$, there exists a constant $L>0$ such that
$\left|g_{i,\mathrm{loc}}^{\prime}(d)\right|\leqslant L\quad\text{for all }d\in\Bigl[\frac{\underline{c}}{2},2\overline{c}\Bigr].$
It then follows that
$|a_{1}-a_{2}|=|g(d_{1})-g(d_{2})|\leqslant L\,|d_{1}-d_{2}|.$ (3.7)
Hence, we have shown that there exists a constant $C>0$ such that
$|a_{1}-a_{2}|\leqslant C\,|d_{1}-d_{2}|,$
which concludes the proof.
Having established Lemma 3.2, we can now complete the proof of the main result of this section.
Proof of Lemma 3.1..
From the equality of pressures we have
$\frac{R_{+}^{\gamma_{+}}}{\alpha_{+}^{\gamma_{+}}}=\frac{R_{-}^{\gamma_{-}}}{(1-\alpha_{+})^{\gamma_{-}}},$
which gives the following equation in $\alpha_{+}$, for $\gamma=\gamma_{+}/\gamma_{-}$:
$\frac{R_{-}}{R_{+}^{\gamma}}\alpha_{+}^{\gamma}+(\alpha_{+}-1)=0.$ (3.8)
Similarly, we obtain for $\alpha_{+}^{\varepsilon}$:
$\frac{R_{-}^{\varepsilon}}{(R_{+}^{\varepsilon})^{\gamma}}(\alpha_{+}^{\varepsilon})^{\gamma}+(\alpha_{+}^{\varepsilon}-1)=0.$ (3.9)
If we denote by $d_{1}=\frac{R_{-}^{\varepsilon}}{(R_{+}^{\varepsilon})^{\gamma}}$ and $d_{2}=\frac{R_{-}}{R_{+}^{\gamma}}$, the hypothesis $d_{2}\in[\underline{c},\overline{c}]$ is verified because $\alpha_{\pm}\in[\underline{a},\overline{a}]\subset(0,1)$, see Remark 1.4, and we can apply Lemma 3.2. It follows that we obtain a unique solution $(\alpha_{+}^{\varepsilon},\alpha_{+})\in(0,1)$ such that equations (3.8) and (3.9) are satisfied and that there exists a constant $C$ such that
$|\alpha_{+}^{\varepsilon}-\alpha_{+}|\leqslant C\left\lvert\frac{R_{-}^{\varepsilon}}{(R_{+}^{\varepsilon})^{\gamma}}-\frac{R_{-}}{R_{+}^{\gamma}}\right\rvert$.
By adding and substracting $\frac{R_{-}}{(R_{+}^{\varepsilon})^{\gamma}}$ and using the assumption $R_{+}^{\varepsilon}>c$, we obtain
$|\alpha_{+}^{\varepsilon}-\alpha_{+}|\leqslant C\left\lvert R_{-}^{\varepsilon}-R_{-}\right\rvert+CR_{-}\left\lvert\frac{1}{(R_{+}^{\varepsilon})^{\gamma}}-\frac{1}{R_{+}^{\gamma}}\right\rvert$.
Finally, using the mean value theorem for the function $f(x)=\frac{1}{x^{\gamma}}$ defined for $x\in\mathbb{R}_{+},x>\underline{x}$, we can conclude the proof.
As a consequence of the previous lemma, we manage to estimate the $L^{2}$ norm of $R_{\pm}^{\varepsilon}-R_{\pm}$ by the entropy and the $L^{1}$ norm of $R_{\pm}^{\varepsilon}-R_{\pm}$, such a result is given in the following proposition.
Proposition 3.3.
Under the assumptions of Theorem 1.6. Let $(\rho^{\varepsilon}_{+},\rho^{\varepsilon}_{-},\alpha^{\varepsilon}_{+},u^{\varepsilon})$ a weak solution of System (LABEL:system_eps), in the sense of Definition 2.1, and $(\rho_{+},\rho_{-},\alpha_{+},u)$ a strong solution of System (1.5), as in Theorem 1.3, there exists $C>0$ depending only on $\underline{a},\overline{a},C_{0}$ such that
$\|R_{\pm}^{\varepsilon}-R_{\pm}\|_{L^{2}}^{2}\leqslant C\varepsilon^{2}{\mathcal{E}}_{1}+C\|R_{+}^{\varepsilon}-R_{+}\|_{L^{1}}+C\|R_{-}^{\varepsilon}-R_{-}\|_{L^{1}}.$
Proof.
We have
$\int_{\Omega}\lvert R_{\pm}^{\varepsilon}-R_{\pm}\rvert^{2}$ $=\int_{\Omega}\lvert\alpha_{\pm}^{\varepsilon}(\rho_{\pm}^{\varepsilon}-\rho_{\pm})+(\alpha_{\pm}^{\varepsilon}-\alpha_{\pm})\rho_{\pm}\rvert^{2}$
$\leqslant C\int_{\Omega}\alpha_{\pm}^{\varepsilon}\lvert\rho_{\pm}^{\varepsilon}-\rho_{\pm}\rvert^{2}+C\int_{\Omega}\rho_{\pm}^{2}\lvert\alpha_{\pm}^{\varepsilon}-\alpha_{\pm}\rvert^{2}.$
The first term is controlled by the relative entropy ${\mathcal{E}}_{1}$ through Inequalities (3.1)–(3.2):
$\int_{\Omega}\alpha_{\pm}^{\varepsilon}\lvert\rho_{\pm}^{\varepsilon}-\rho_{\pm}\rvert^{2}$ $=\int_{\{\lvert\rho_{\pm}^{\varepsilon}-\rho_{\pm}\rvert\leqslant 1\}}\alpha_{\pm}^{\varepsilon}\lvert\rho_{\pm}^{\varepsilon}-\rho_{\pm}\rvert^{2}+\int_{\{\lvert\rho_{\pm}^{\varepsilon}-\rho_{\pm}\rvert>1\}}\alpha_{\pm}^{\varepsilon}\lvert\rho_{\pm}^{\varepsilon}-\rho_{\pm}\rvert^{\gamma_{\pm}}$
$\leqslant C\varepsilon^{2}{\mathcal{E}}_{1}.$ (3.10)
where we have used $\gamma_{+},\gamma_{-}\geqslant 2$.
For the second term, we want to show that
$\int_{\Omega}\rho_{\pm}^{2}\lvert\alpha_{\pm}^{\varepsilon}-\alpha_{\pm}\rvert^{2}\leqslant C\varepsilon^{2}{\mathcal{E}}_{1}+C\sum_{i=+,-}\int_{\Omega}\lvert R_{i}^{\varepsilon}-R_{i}\rvert$. (3.11)
Let us consider $\tilde{C_{0}}>0$, depending only on $\overline{a}$ and $C_{0}$ such that $R_{\pm}\leqslant\tilde{C_{0}}$.
  • In the set $\Omega_{1,\pm}:=\{x,\ (R_{\pm}^{\varepsilon}+R_{\pm})\leqslant 2\tilde{C}_{0}\}$, we use the equality:
    $\rho_{\pm}^{2}\left(\alpha_{\pm}^{\varepsilon}-\alpha_{\pm}\right)^{2}$ $=(\alpha_{\pm}^{\varepsilon})^{2}\left(\rho_{\pm}-\rho_{\pm}^{\varepsilon}\right)^{2}+2\alpha_{\pm}^{\varepsilon}\rho_{\pm}\left(R_{\pm}^{\varepsilon}-R_{\pm}\right)+\left(R_{\pm}^{2}-(R_{\pm}^{\varepsilon})^{2}\right).$
    For the last term, we use $(R_{\pm}^{\varepsilon}+R_{\pm})\leqslant 2\tilde{C}_{0}$ to get:
    $\int_{\Omega_{1,\pm}}\lvert(R_{\pm}^{\varepsilon})^{2}-R_{\pm}^{2}\rvert\leqslant\int_{\Omega_{1,\pm}}\lvert R_{\pm}^{\varepsilon}-R_{\pm}\rvert(R_{\pm}^{\varepsilon}+R_{\pm})\leqslant C\int_{\Omega}\lvert R_{\pm}^{\varepsilon}-R_{\pm}\rvert$.
    Therefore in the first set we obtain
    $\int_{\Omega_{1,\pm}}\rho_{\pm}^{2}\lvert\alpha_{\pm}^{\varepsilon}-\alpha_{\pm}\rvert^{2}$ $\leqslant\int_{\Omega}\alpha_{\pm}^{\varepsilon}\lvert\rho_{\pm}^{\varepsilon}-\rho_{\pm}\rvert^{2}+2\lVert\rho_{\pm}\rVert_{L^{\infty}}\int_{\Omega}\lvert R_{\pm}^{\varepsilon}-R_{\pm}\rvert+C\int_{\Omega}\lvert R_{\pm}^{\varepsilon}-R_{\pm}\rvert$
    $\leqslant\int_{\Omega}\alpha_{\pm}^{\varepsilon}\lvert\rho_{\pm}^{\varepsilon}-\rho_{\pm}\rvert^{2}+C\int_{\Omega}\lvert R_{\pm}^{\varepsilon}-R_{\pm}\rvert$,
    which, by using (3.10), can be estimated by the right-hand side terms of (3.11).
  • Now in the complementary set $\Omega_{2,\pm}:=\{x,\ (R_{\pm}^{\varepsilon}+R_{\pm})>2\tilde{C}_{0}\}$, we have $R_{\pm}^{\varepsilon}>\tilde{C}_{0}$ and the result comes directly from Lemma 3.1:
[…]
This ends the proof of (3.11) and of this proposition. ∎
3.2 Csiszár-Kullback-Pinsker-type inequality
We will now present an inequality close to the Csiszar-Kullback-Pinsker inequality.
Proposition 3.4.
Let $f$ a nonnegative and $g$ a positive real functions in $L^{1}$ such that $\int f=\int g$. Then
$\|f-g\|_{L^{1}}\leqslant\int f\ln_{+}\frac{f}{g}$
and
$\int f\Big|\ln\frac{f}{g}\Big|\leqslant 2\int f\ln_{+}\frac{f}{g}$
Proof.
As
$\int_{\{g<f\}}f-g=-\int_{\{g\geqslant f\}}f-g$
we have
$\|f-g\|_{L^{1}}=\int_{\{g<f\}}f-g+\int_{\{g\geqslant f\}}g-f=2\int_{\{g<f\}}f-g=2\int_{\{g<f\}}g\Big(\frac{f}{g}-1\Big).$
By convexity of $h(r)=r\ln r$ we have $r-1\leqslant r\ln r$ for all $r>0$, hence
$\|f-g\|_{L^{1}}\leqslant 2\int_{\{g<f\}}g\frac{f}{g}\ln\frac{f}{g}=2\int f\ln_{+}\frac{f}{g}$
which gives the first result.
For the second inequality, we compute
$\int_{\{g\geqslant f\}}f\ln\frac{f}{g}=\int_{\{g\geqslant f>0\}}f(\ln g-\ln f)\leqslant\int_{\{g\geqslant f>0\}}f\frac{1}{f}(g-f)\leqslant\|f-g\|_{L^{1}}$
and we end this proof by applying the first inequality. ∎
Remark 3.5.
The classical Csiszár–Kullback–Pinsker inequality provides a control of the squared $L^{1}$-distance by the relative entropy with logarithmic structure (the Kullback–Leibler divergence). However in our setting, it is essential to control the $L^{1}$-norm itself rather than its square because an estimate involving $\|f-g\|_{L^{1}}^{2}$ would not be sufficient for our purposes, as it would not yield a differential inequality compatible with a direct application of Gronwall’s lemma. For this reason, we establish a refined inequality that provides direct control of $\|f-g\|_{L^{1}}$. The price to pay is a weaker logarithmic structure in the relative entropy component: the natural logarithm is replaced by its positive part $\ln_{+}$.
We will apply the previous proposition for $R_{\pm}^{\varepsilon}$ and $R_{\pm}$, where $R_{\pm}^{\varepsilon},R_{\pm}\geqslant 0$ and $R_{\pm}>\underline{C}$.
Corollary 3.6.
If $R_{\pm}^{\varepsilon}$ and $R_{\pm}$ are nonnegative real functions in $L^{1}(\Omega)$ such that $\int_{\Omega}R_{\pm}^{\varepsilon}=\int_{\Omega}R_{\pm}$. Then
$\|R_{\pm}^{\varepsilon}-R_{\pm}\|_{L^{1}}\leqslant\int_{\Omega}R_{\pm}^{\varepsilon}\ln_{+}\frac{R_{\pm}^{\varepsilon}}{R_{\pm}}.$
4 Strong convergence of the partial densities $R_{\pm}^{\varepsilon}$
In this section, we establish the strong convergence of the partial densities $R_{\pm}^{\varepsilon}$, relying on the relative entropy framework and the structural estimates derived in the previous sections.
From (1.7) and the energy inequality (2.4), we first deduce uniform-in-$\varepsilon$ bounds for the following quantities:
  • The kinetic energy weighted by the total density,
    $\big(\alpha_{+}^{\varepsilon}\rho_{+}^{\varepsilon}+\alpha_{-}^{\varepsilon}\rho_{-}^{\varepsilon}\big)|u^{\varepsilon}|^{2}\in L^{\infty}(0,\infty;L^{1}(\Omega)),$
  • The rescaled internal energies,
    $\frac{1}{\varepsilon^{2}}\alpha_{i}^{\varepsilon}(\rho_{i}^{\varepsilon})^{\gamma_{i}}\in L^{\infty}(0,\infty;L^{1}(\Omega)),\quad i=+,-,$
  • The velocity gradients,
    $\nabla u^{\varepsilon},\;\operatorname{div}u^{\varepsilon}\in L^{2}(0,\infty;L^{2}(\Omega)).$
In particular, this implies that the partial mass densities $R_{i}^{\varepsilon}=\alpha_{i}^{\varepsilon}\rho_{i}^{\varepsilon}$ are uniformly bounded in $L^{\infty}(0,\infty;L^{\gamma_{i}}(\Omega))$ for $i=+,$-, and that the velocity field $u^{\varepsilon}$ is uniformly bounded in $L^{2}(0,T;H_{0}^{1}(\Omega))$ for all $T>0$ in the case of Dirichlet boundary conditions, by the Poincaré inequality.
4.1 Uniform bound for ${\mathcal{E}}_{2}$.
Recall that we have defined the following entropy in (2.3)
$\begin{split}{\mathcal{E}}_{2}\Big(\rho^{\varepsilon}_{+},\rho^{\varepsilon}_{-},\alpha^{\varepsilon}_{+},u^{\varepsilon}|\rho_{+},\rho_{-},\alpha_{+},u\Big)=\int_{\Omega}R_{+}^{\varepsilon}\ln_{+}\left(\frac{R_{+}^{\varepsilon}}{R_{+}}\right)+\int_{\Omega}R_{-}^{\varepsilon}\ln_{+}\left(\frac{R_{-}^{\varepsilon}}{R_{-}}\right),\end{split}$
and from Corollary 2.6
${\mathcal{E}}_{2}(t)-{\mathcal{E}}_{2}(0)\leqslant C\sum_{i=+,-}\int_{0}^{t}\|R_{i}^{\varepsilon}-R_{i}\|_{L^{2}(\Omega)}\Big(\|\nabla(u^{\varepsilon}-u)\|_{L^{2}(\Omega)}+\|\operatorname{div}(u-u^{\varepsilon})\|_{L^{2}(\Omega)}\Big)+C.$
Applying Young’s inequality with an arbitrary $\delta_{1}>0$ we obtain:
$\|R_{i}^{\varepsilon}-R_{i}\|_{L^{2}(\Omega)}\|\nabla(u^{\varepsilon}-u)\|_{L^{2}(\Omega)}$ $\leqslant\frac{1}{2\delta_{1}}\|R_{\pm}^{\varepsilon}-R_{\pm}\|_{L^{2}}^{2}+\frac{\delta_{1}}{2}\lVert\nabla(u-u^{\varepsilon})\rVert_{L^{2}(\Omega)}^{2}.$
Then using Proposition 3.3:
$\|R_{i}^{\varepsilon}-R_{i}\|_{L^{2}(\Omega)}\|\nabla(u^{\varepsilon}-u)\|_{L^{2}(\Omega)}$ $\leqslant\frac{C}{\delta_{1}}\left(\varepsilon^{2}{\mathcal{E}}_{1}+\sum_{i=+,-}\|R_{i}^{\varepsilon}-R_{i}\|_{L^{1}}\right)+\frac{\delta_{1}}{2}\lVert\nabla(u-u^{\varepsilon})\rVert_{L^{2}(\Omega)}^{2}.$
And similarly we get for the second term
$\|R_{i}^{\varepsilon}-R_{i}\|_{L^{2}(\Omega)}\|\operatorname{div}(u-u^{\varepsilon})\|_{L^{2}(\Omega)}$ $\leqslant\frac{C}{\delta_{2}}\left(\varepsilon^{2}{\mathcal{E}}_{1}+\sum_{i=+,-}\|R_{i}^{\varepsilon}-R_{i}\|_{L^{1}}\right)+\frac{\delta_{2}}{2}\|\operatorname{div}(u-u^{\varepsilon})\|^{2}_{L^{2}(\Omega)}.$
Therefore, using Corollary 3.6,
${\mathcal{E}}_{2}(t)-{\mathcal{E}}_{2}(0)\leqslant C(\delta_{1},\delta_{2})\int_{0}^{t}(\varepsilon^{2}{\mathcal{E}}_{1}+{\mathcal{E}}_{2})+\frac{\delta_{2}}{2}\int_{0}^{t}\|\operatorname{div}(u-u^{\varepsilon})\|^{2}_{L^{2}(\Omega)}\\ +\frac{\delta_{1}}{2}\int_{0}^{t}\lVert\nabla(u-u^{\varepsilon})\rVert_{L^{2}(\Omega)}^{2}+C.$ (4.1)
4.2 Uniform bound for ${\mathcal{E}}_{1}$.
Recall that thanks to Propositions 2.3 and 2.4, we have
${\mathcal{E}}_{1}(t)-{\mathcal{E}}_{1}(0)$ $+\mu\int_{0}^{t}\int_{\Omega}\left(\nabla(u^{\varepsilon}-u)\right)^{2}+(\lambda+\mu)\int_{0}^{t}\int_{\Omega}\left(\operatorname{div}(u^{\varepsilon}-u)\right)^{2}$ (4.2)
$\leqslant$ $\mu\int_{0}^{t}\int_{\Omega}\frac{\rho^{\varepsilon}-\rho}{\rho}\Delta u(u-u^{\varepsilon})$ (4.3)
$+\int_{0}^{t}\int_{\Omega}\rho^{\varepsilon}(u-u^{\varepsilon})\cdot\nabla u\cdot(u^{\varepsilon}-u)$ (4.4)
$-\int_{0}^{t}\int_{\Omega}\frac{\rho^{\varepsilon}}{\rho}\nabla\Pi(u-u^{\varepsilon}).$ (4.5)
We deduce an estimate for the term (4.4)
$\left\lvert\int_{\Omega}\rho^{\varepsilon}(u-u^{\varepsilon})\cdot\nabla u\cdot(u^{\varepsilon}-u)\right\rvert$ $\leqslant c\lVert\nabla u\rVert_{L^{\infty}}\int_{\Omega}\rho^{\varepsilon}\left\lvert u^{\varepsilon}-u\right\rvert^{2}$
$\leqslant c\lVert\nabla u\rVert_{L^{\infty}}{\mathcal{E}}_{1}\Big(\rho^{\varepsilon}_{+},\rho^{\varepsilon}_{-},\alpha^{\varepsilon}_{+},u^{\varepsilon}|\rho_{+},\rho_{-},\alpha_{+},u\Big).$
Now we focus on the term (4.3). Recalling that $\rho^{\varepsilon}-\rho=\sum_{i=+,-}R_{i}^{\varepsilon}-R_{i}$, we use Hölder inequality ($1/2+1/3+1/6=1$), to get
$\left\lvert\int_{\Omega}\frac{R_{\pm}^{\varepsilon}-R_{\pm}}{\rho}\Delta u(u-u^{\varepsilon})\right\rvert$ $\leqslant\lVert\frac{1}{\rho}\Delta u\rVert_{L^{3}(\Omega)}\lVert u-u^{\varepsilon}\rVert_{L^{6}(\Omega)}\left(\int_{\Omega}\lvert R_{\pm}^{\varepsilon}-R_{\pm}\rvert^{2}\right)^{1/2}.$
Therefore, using Young’s inequality with an arbitrary $\delta>0$,
$\left\lvert\mu\int_{\Omega}\frac{\alpha_{\pm}^{\varepsilon}\rho_{\pm}^{\varepsilon}-\alpha_{\pm}\rho_{\pm}}{\rho}\Delta u(u-u^{\varepsilon})\right\rvert$ $\leqslant\frac{C}{\delta}\lVert\frac{1}{\rho}\Delta u\rVert_{L^{3}(\Omega)}^{2}\int_{\Omega}\lvert R_{\pm}^{\varepsilon}-R_{\pm}\rvert^{2}+C\delta\lVert u-u^{\varepsilon}\rVert_{L^{6}(\Omega)}^{2}.$
Using Sobolev injections and Poincaré inequality for the second term
$\lVert u-u^{\varepsilon}\rVert_{L^{6}(\Omega)}^{2}\leqslant C\lVert\nabla(u-u^{\varepsilon})\rVert^{2}_{L^{2}(\Omega)}.$
Finally, using Proposition 3.3 and Corollary 3.6, we obtain
$\left\lvert\mu\int_{\Omega}\frac{\alpha_{\pm}^{\varepsilon}\rho_{\pm}^{\varepsilon}-\alpha_{\pm}\rho_{\pm}}{\rho}\Delta u(u-u^{\varepsilon})\right\rvert$ $\leqslant\frac{C}{\delta}\varepsilon^{2}{\mathcal{E}}_{1}+\frac{C}{\delta}{\mathcal{E}}_{2}+\delta C\lVert\nabla(u-u^{\varepsilon})\rVert^{2}_{L^{2}(\Omega)}.$
We end by finding an estimate for the term (4.5). We have
$\left\lvert\int_{\Omega}-\frac{\rho^{\varepsilon}}{\rho}\nabla\Pi(u-u^{\varepsilon})\right\rvert$ $=\left\lvert\int_{\Omega}\frac{\rho-\rho^{\varepsilon}}{\rho}\nabla\Pi(u-u^{\varepsilon})-\nabla\Pi(u-u^{\varepsilon})\right\rvert$
$\leqslant\int_{\Omega}\left\lvert\frac{\rho-\rho^{\varepsilon}}{\rho}\nabla\Pi(u-u^{\varepsilon})\right\rvert+\int_{\Omega}\left\lvert\nabla\Pi(u-u^{\varepsilon})\right\rvert$.
For the first term we can use the same method as we just did previously to find an estimate for (4.3), where now $\nabla\Pi$ replaces $\Delta u$. We have
$\int_{\Omega}\left\lvert\frac{\rho-\rho^{\varepsilon}}{\rho}\nabla\Pi(u-u^{\varepsilon})\right\rvert\leqslant\frac{C}{\delta}\varepsilon^{2}{\mathcal{E}}_{1}+\frac{C}{\delta}{\mathcal{E}}_{2}+\delta C\lVert\nabla(u-u^{\varepsilon})\rVert^{2}_{L^{2}(\Omega)}.$
Then for the second term we just use the Poincaré inequality:
$\int_{\Omega}\left\lvert\nabla\Pi(u-u^{\varepsilon})\right\rvert\leqslant\frac{1}{\delta}\lVert\nabla\Pi\rVert_{L^{2}}^{2}+\delta C\lVert\nabla(u^{\varepsilon}-u)\rVert_{L^{2}}^{2}.$
Therefore, we have the following inequality for ${\mathcal{E}}_{1}$:
${\mathcal{E}}_{1}(t)-{\mathcal{E}}_{1}(0)$ $+\mu\int_{0}^{t}\int_{\Omega}\left\lvert\nabla(u^{\varepsilon}-u)\right\rvert^{2}+(\lambda+\mu)\int_{0}^{t}\int_{\Omega}\left\lvert\operatorname{div}(u^{\varepsilon}-u)\right\rvert^{2}$
$\leqslant C\int_{0}^{t}({\mathcal{E}}_{1}+{\mathcal{E}}_{2})+C\delta\int_{0}^{t}\lVert\nabla(u^{\varepsilon}-u)\rVert_{L^{2}}^{2}+\frac{1}{\delta}\int_{0}^{t}\lVert\nabla\Pi\rVert_{L^{2}}^{2}.$ (4.6)
4.3 Uniform bound for $\mathcal{E}$}
Putting together the estimates for ${\mathcal{E}}_{2}$ and ${\mathcal{E}}_{1}$, see (4.1) and (4.2), we conclude that
${\mathcal{E}}(t)-{\mathcal{E}}(0)$ $+\mu\int_{0}^{t}\int_{\Omega}\left\lvert\nabla(u^{\varepsilon}-u)\right\rvert^{2}+(\lambda+\mu)\int_{0}^{t}\int_{\Omega}\left\lvert\operatorname{div}(u^{\varepsilon}-u)\right\rvert^{2}$
$\leqslant C(\delta_{0},\delta_{2})\int_{0}^{t}{\mathcal{E}}+\delta_{0}\int_{0}^{t}\lVert\nabla(u^{\varepsilon}-u)\rVert_{L^{2}}^{2}$ (4.7)
$\qquad+\delta_{2}\int_{0}^{t}\|\operatorname{div}(u-u^{\varepsilon})\|^{2}_{L^{2}(\Omega)}+C(\delta_{0})\int_{0}^{t}\lVert\nabla\Pi\rVert_{L^{2}}^{2}.$
Setting $\delta_{0}=\mu/2$ and $\delta_{2}=(\lambda+\mu)/2$, a Gronwall-type argument allows us to state that ${\mathcal{E}}_{1}(t)$ and ${\mathcal{E}}_{2}(t)$ are uniformly bounded on $[0,T]$.
4.4 Convergence of the densities and volume fractions
As all the problem is symmetric in terms of $\gamma_{+}$, $\gamma_{-}$, we can assume without loss of generality that $\gamma_{+}>\gamma_{-}$. Alternatively, we can exchange $\alpha_{\pm},\rho_{\pm}$ by $\alpha_{\mp},\rho_{\mp}$ in this subsection. Hence $\gamma=\gamma_{+}/\gamma_{-}>1$. The equality of pressures gives
$\rho_{-}=\rho_{+}^{\gamma},$
so that we can write
$\mathcal{E}_{1}$ $\geqslant\int_{\Omega}\frac{\alpha_{+}^{\varepsilon}}{\varepsilon^{2}(\gamma_{+}-1)}\lvert\rho_{+}^{\varepsilon}-\rho_{+}\rvert^{2}+\int_{\Omega}\frac{\alpha_{-}^{\varepsilon}}{\varepsilon^{2}(\gamma_{-}-1)}\lvert\rho_{-}^{\varepsilon}-\rho_{-}\rvert^{2}$
$\geqslant\int_{\Omega}\frac{\alpha_{+}^{\varepsilon}}{\varepsilon^{2}(\gamma_{+}-1)}\lvert\rho_{+}^{\varepsilon}-\rho_{+}\rvert^{2}+\int_{\Omega}\frac{\alpha_{-}^{\varepsilon}}{\varepsilon^{2}(\gamma_{-}-1)}\lvert(\rho_{+}^{\varepsilon})^{\gamma}-\rho_{+}^{\gamma}\rvert^{2}.$
For all $x\in\Omega$ such that $\rho_{+}^{\varepsilon}<\rho_{+}/2$, we have
$|\rho_{+}^{\varepsilon\gamma}-\rho_{+}^{\gamma}|=\rho_{+}^{\gamma}-\rho_{+}^{\varepsilon\gamma}\geqslant\rho_{+}^{\gamma}-(\rho_{+}/2)^{\gamma}\geqslant(1-2^{-\gamma})\frac{\rho_{+}^{\gamma}}{\rho_{+}}|\rho_{+}^{\varepsilon}-\rho_{+}|.$
because $|\rho_{+}^{\varepsilon}-\rho_{+}|=\rho_{+}-\rho_{+}^{\varepsilon}\leqslant\rho_{+}$
Otherwise, for all $x$ such that $\rho_{+}^{\varepsilon}>\rho_{+}/2$, we have by the mean value theorem applied to $F(x)=x^{1/\gamma}$
$|\rho_{+}^{\varepsilon}-\rho_{+}|\leqslant\sup_{x>(\rho_{+}/2)^{1/\gamma}}F^{\prime}(x)|\rho_{+}^{\varepsilon\gamma}-\rho_{+}^{\gamma}|=\frac{1}{\gamma}(\rho_{+}/2)^{\frac{1}{\gamma}-1}|\rho_{+}^{\varepsilon\gamma}-\rho_{+}^{\gamma}|$
Hence, we obtain the lower bound
$\mathcal{E}_{1}$ $\geqslant C\int_{\Omega}\frac{\alpha_{+}^{\varepsilon}}{\varepsilon^{2}(\gamma_{+}-1)}\lvert\rho_{+}^{\varepsilon}-\rho_{+}\rvert^{2}+\int_{\Omega}\frac{\alpha_{-}^{\varepsilon}}{\varepsilon^{2}(\gamma_{-}-1)}\lvert\rho_{+}^{\varepsilon}-\rho_{+}\rvert^{2}.$
for some constant $C>0$.
Since $\mathcal{E}_{1}$ is uniformly bounded, we deduce that
$\rho_{+}^{\varepsilon}\to\rho_{+}\quad\text{in }L^{\infty}((0,T);L^{2}(\Omega))\quad\text{as }\varepsilon\to 0.$
Doing a similar type of analysis using the equality of pressures, we also obtain
$\rho_{-}^{\varepsilon}\to\rho_{-}\quad\text{in }L^{\infty}((0,T);L^{2}(\Omega)).$
Furthermore, since $\alpha_{\pm}^{\varepsilon}\leqslant 1$, the sequences $(\alpha_{\pm}^{\varepsilon})_{\varepsilon}$ are uniformly bounded in $L^{1}(\Omega)$ and $L^{2}(\Omega)$. Hence, up to extraction of a subsequence (still denoted by $\alpha_{\pm}^{\varepsilon}$), there exist limits $\bar{\alpha}_{\pm}$ such that
$\alpha_{\pm}^{\varepsilon}\rightharpoonup\bar{\alpha}_{\pm}\quad\text{weakly-$*$ in }L^{\infty}((0,T);L^{2}(\Omega)),$
with
$\bar{\alpha}_{+}+\bar{\alpha}_{-}=1.$
Finally, using the product of a strongly convergent sequence and a weakly convergent sequence, we deduce that, for $i=+,$-,
$R_{i}^{\varepsilon}=\alpha_{i}^{\varepsilon}\rho_{i}^{\varepsilon}\rightharpoonup\bar{\alpha}_{i}\rho_{i}\quad\text{weakly in }L^{2}(0,T;L^{2}(\Omega)).$
4.5 Strong convergence of the partial densities $R_{\pm}^{\varepsilon}$ and weak convergence of $\operatorname{div}(u^{\varepsilon})$
We recall that $u^{\varepsilon}$ is uniformly bounded in $L^{2}(0,T;H^{1}(\Omega))$ by the energy estimate. By Sobolev embeddings in space dimension $d=3$, we have for $p\in[2,6]$
$H^{1}(\Omega)\hookrightarrow L^{p}(\Omega),$
so that, taking $p=6$, $u^{\varepsilon}$ is uniformly bounded in $L^{2}(0,T;L^{6}(\Omega))$.
Since $R_{\pm}^{\varepsilon}\in L^{\infty}(0,T;L^{2}(\Omega))$, Hölder’s inequality implies that $R_{\pm}^{\varepsilon}u^{\varepsilon}$ is uniformly bounded in $L^{2}(0,T;L^{3/2}(\Omega))$.
Consequently, $\operatorname{div}(R_{\pm}^{\varepsilon}u^{\varepsilon})$ is uniformly bounded in $L^{2}(0,T;W^{-1,3/2}(\Omega))$ in the sense of distributions.
Using the mass conservation equations, we deduce that $\partial_{t}R_{\pm}^{\varepsilon}$ is uniformly bounded in $L^{2}(0,T;W^{-1,3/2}(\Omega))$. Applying a version of the Aubin–Lions lemma together with the compact and continuous embeddings
$L^{2}(\Omega)\hookrightarrow\hookrightarrow H^{-1}(\Omega)\hookrightarrow W^{-1,\frac{3}{2}}(\Omega),$
we infer that, up to the extraction of a subsequence,
$R_{\pm}^{\varepsilon}\to\bar{R}_{\pm}\quad\text{strongly in }L^{2}(0,T;H^{-1}(\Omega)).$
By weak strong uniqueness of the limit, we conclude that
$\bar{R}_{\pm}=\bar{\alpha}_{\pm}\rho_{\pm}.$
Moreover, since $\mathcal{E}_{1}$ is uniformly bounded and $R_{\pm}^{\varepsilon}$ is bounded in $L^{\infty}(0,T;L^{2}(\Omega))$, Poincaré’s inequality ensures that the quantity
$R_{\pm}^{\varepsilon}|u^{\varepsilon}-u|^{2}$
is uniformly bounded in $L^{1}((0,T)\times\Omega)$. In view of the uniform bounds on $R_{\pm}^{\varepsilon}$, we therefore infer that the sequence $(u^{\varepsilon}-u)_{\varepsilon}$ is uniformly bounded in $L^{2}((0,T)\times\Omega)$. Since $L^{2}((0,T)\times\Omega)$ is reflexive, there exists a function $v\in L^{2}((0,T)\times\Omega)$ and a subsequence (not relabeled) such that
$u^{\varepsilon}-u\rightharpoonup v\quad\text{weakly in }L^{2}((0,T)\times\Omega).$
Passing to the limit in the mass equations, we obtain, for $i=+,$-,
$\partial_{t}(\bar{\alpha}_{i}\rho_{i})+\operatorname{div}\big(\bar{\alpha}_{i}\rho_{i}(u+v)\big)=0,$
which yields
$\partial_{t}\bar{\alpha}_{i}+\operatorname{div}\big(\bar{\alpha}_{i}(u+v)\big)=0.$
Summing the two equations for $i=+,$-, and using $\bar{\alpha}_{+}+\bar{\alpha}_{-}=1$ together with $\operatorname{div}u=0$, we deduce
$\operatorname{div}v=0.$
Therefore, we conclude that
$\operatorname{div}u^{\varepsilon}\rightharpoonup 0\quad\text{weakly in }L^{2}((0,T)\times\Omega).$ (4.8)
Using a standard compactness argument for Navier Stokes, we also get a strong convergence of $\rho^{\varepsilon}u^{\varepsilon}$ in $L^{2}(0,T;H^{-1}(\Omega))$, which allows to pass at the limit in the momentum equation for any test function divergence free. This allows to prove that $(\overline{R}_{+}+\overline{R}_{-},u+v)$ satisfies the limit system (1.5) in the weak sense. By weak strong uniqueness, we have that $\overline{R}_{+}+\overline{R}_{-}=\rho$ and $v=0$. By Remark 1.4 this means that $\overline{\alpha}_{\pm}=\alpha_{\pm}$ and $\overline{R}_{\pm}=R_{\pm}$.
5 End of proof
We will now find a new estimate for ${\mathcal{E}}_{1}$ and use the convergences found above to conclude.
We first find a new bound for the term (4.3). Recalling that $\rho^{\varepsilon}-\rho=\sum_{i=+,-}R_{i}^{\varepsilon}-R_{i}$, we use Cauchy-Schwartz in dual spaces $H^{1}(\Omega)$ and $H^{-1}(\Omega)$, to get
$\left\lvert\int_{\Omega}\frac{R_{\pm}^{\varepsilon}-R_{\pm}}{\rho}\Delta u(u-u^{\varepsilon})\right\rvert$ $\leqslant\left\lVert\frac{1}{\rho}\Delta u(u-u^{\varepsilon})\right\rVert_{H^{1}(\Omega)}\lVert R_{\pm}^{\varepsilon}-R_{\pm}\rVert_{H^{-1}(\Omega)}$
$\leqslant C\lVert\Delta u\rVert_{H^{1}(\Omega)}\lVert u-u^{\varepsilon}\rVert_{H^{1}(\Omega)}\lvert R_{\pm}^{\varepsilon}-R_{\pm}\rVert_{H^{-1}(\Omega)}.$
Then, to find a new estimate for the term (4.5), we write
$\left\lvert\int_{0}^{t}\int_{\Omega}-\frac{\rho^{\varepsilon}}{\rho}\nabla\Pi(u-u^{\varepsilon})\right\rvert$ $\leqslant\int_{0}^{t}\left\lvert\int_{\Omega}\frac{\rho-\rho^{\varepsilon}}{\rho}\nabla\Pi(u-u^{\varepsilon})\right\rvert+\left\lvert\int_{0}^{t}\int_{\Omega}\nabla\Pi(u-u^{\varepsilon})\right\rvert$.
and similarly, we get for the first term
$\left\lvert\int_{\Omega}\frac{\rho-\rho^{\varepsilon}}{\rho}\nabla\Pi(u-u^{\varepsilon})\right\rvert\leqslant C\lVert\nabla\Pi\rVert_{H^{1}(\Omega)}\lVert u-u^{\varepsilon}\rVert_{H^{1}(\Omega)}\lvert R_{\pm}^{\varepsilon}-R_{\pm}\rVert_{H^{-1}(\Omega)}.$
Therefore, we deduce
$\mathcal{E}_{1}(t)-\mathcal{E}_{1}(0)+\mu\int_{0}^{t}\int_{\Omega}\lvert\nabla(u^{\varepsilon}-u)\rvert^{2}-(\lambda+\mu)\int_{0}^{t}\int_{\Omega}\lvert\operatorname{div}(u^{\varepsilon}-u)\rvert^{2}$
$\quad\leqslant C\int_{0}^{t}\mathcal{E}_{1}(s)\,ds+C\int_{0}^{t}\lVert u-u^{\varepsilon}\rVert_{H^{1}(\Omega)}\lvert R_{\pm}^{\varepsilon}-R_{\pm}\rVert_{H^{-1}(\Omega)}\,ds+\Big\lvert\int_{0}^{t}\int_{\Omega}\Pi\operatorname{div}(u-u^{\varepsilon})\,dx\,ds\Big\rvert$.
Applying Gronwall’s lemma, we finally obtain
$\mathcal{E}_{1}(t)\leqslant e^{Ct}\left[\mathcal{E}_{1}(0)+C\int_{0}^{t}\lVert u-u^{\varepsilon}\rVert_{H^{1}(\Omega)}\lvert R_{\pm}^{\varepsilon}-R_{\pm}\rVert_{H^{-1}(\Omega)}\,ds+\Big\lvert\int_{0}^{t}\int_{\Omega}\Pi\operatorname{div}(u-u^{\varepsilon})\,dx\,ds\Big\rvert\right$].
Assuming that the initial energies satisfy
$\mathcal{E}_{1}(0)\longrightarrow 0\quad\text{as }\varepsilon\to 0,$
and recalling that
$R_{\pm}^{\varepsilon}\to R_{\pm}\quad\text{strongly in }L^{2}(0,T;H^{-1}(\Omega)).$
and
$\operatorname{div}u^{\varepsilon}\rightharpoonup 0\quad\text{in }L^{2}((0,T)\times\Omega),$
we immediately deduce from the previously established Gronwall estimate that
$\mathcal{E}_{1}(t)\longrightarrow 0\quad\text{as }\varepsilon\to 0.$
Combining the energy estimates and the previously established bounds, we conclude the convergence of the approximate solutions $(\alpha_{\pm}^{\varepsilon},\rho_{\pm}^{\varepsilon},u^{\varepsilon})$ as $\varepsilon\to 0$.
From the Gronwall estimate and the hypothesis that the initial energies satisfy
$\mathcal{E}_{1}(0)\longrightarrow 0\quad\text{as }\varepsilon\to 0,$
we deduce that the velocity converges strongly:
$u^{\varepsilon}\longrightarrow u\quad\text{strongly in }L^{2}(0,T;L^{2}(\Omega)).$
The gradients $(\nabla u^{\varepsilon})_{\varepsilon}$ are uniformly bounded in $L^{2}$, so that up to extraction,
$\nabla u^{\varepsilon}\rightharpoonup\nabla u\quad\text{weakly in }L^{2}((0,T)\times\Omega),$
and in particular, the divergence satisfies
$\operatorname{div}u^{\varepsilon}\rightharpoonup 0\quad\text{weakly in }L^{2}((0,T)\times\Omega),$
which shows that the limit velocity $u$ is incompressible.
In summary, the sequence of approximate solutions converges to the expected limit
$(\alpha_{+},\alpha_{-},\rho_{+},\rho_{-},u)$
with the convergence properties
$u^{\varepsilon}$ $\longrightarrow u$ $\text{strongly in }L^{2}(0,T;L^{2}(\Omega)),$
$\nabla u^{\varepsilon}$ $\rightharpoonup\nabla u$ $\text{weakly in }L^{2}((0,T)\times\Omega),$
$\operatorname{div}u^{\varepsilon}$ $\rightharpoonup 0$ $\text{weakly in }L^{2}((0,T)\times\Omega),$
$\rho_{i}^{\varepsilon}$ $\longrightarrow\rho_{i}$ $\text{strongly in }L^{\infty}(0,T;L^{2}(\Omega)),$
$\alpha_{i}^{\varepsilon}$ $\longrightarrow\alpha_{i}$ $\text{weak-$*$ in }L^{\infty}(0,T;L^{2}(\Omega)),$
$R_{i}^{\varepsilon}$ $\longrightarrow R_{i}$ $\text{strongly in }L^{2}(0,T;H^{-1}(\Omega)),\quad i=+,-.$
This convergence framework is standard in the context of singular limits for compressible Navier–Stokes equations and allows the identification of the limit solution as the physically relevant incompressible two-phase flow.
Acknowledgements. The author would like to thank Didier Bresch and Christophe Lacave for their many tips on writing this text. She also gratefully acknowledges the University Savoie Mont Blanc for her PhD position fellowship linked to the CPJ managed by Christophe Lacave. She also wants to thank for some support of the ANR under France 2030 bearing the reference ANR-23-EXMA-004 (Complexflows project) related to the PEPR Maths-ViVEs. The author wants also to gratefully acknowledge the partial support by the Agence Nationale pour la Recherche grant ANR-23-CE40-0014-01 (ANR Bourgeons).